### v1

# Light single-gluon hybrid states with various (exotic) quantum numbers

Abstract
We apply the QCD sum rule method to study the light single-gluon hybrid states with various (exotic) quantum numbers. We construct twenty-four single-gluon hybrid currents, and use eighteen of them to calculate the masses of forty-four single-gluon hybrid states with the quark-gluon contents $\bar{q}qg$ ($q=u/d$) and $\bar{s}sg$. We concentrate on the hybrid states with the exotic quantum number $J^{PC}=1^{-+}$, whose masses and widths are calculated to be $M_{|\bar{q}qg;1^{-}1^{-+}\rangle}=1.67^{+0.15}_{-0.17}$ GeV, $\Gamma_{|\bar{q}qg;1^{-}1^{-+}\rangle}=530^{+540}_{-330}$ MeV, $M_{|\bar{q}qg;0^{+}1^{-+}\rangle}=1.67^{+0.15}_{-0.17}$ GeV, $\Gamma_{|\bar{q}qg;0^{+}1^{-+}\rangle}=120^{+160}_{-110}$ MeV, $M_{|\bar{s}sg;0^{+}1^{-+}\rangle}=1.84^{+0.14}_{-0.15}$ GeV, and $\Gamma_{|\bar{s}sg;0^{+}1^{-+}\rangle}=100^{+110}_{-~{}80}$ MeV. Our results support the interpretations of the $\pi_{1}(1600)$ and $\eta_{1}(1855)$ as the hybrid states $|\bar{q}qg;1^{-}1^{-+}\rangle$ and $|\bar{s}sg;0^{+}1^{-+}\rangle$, respectively. Considering the uncertainties, our results suggest that the $\pi_{1}(1600)$ and $\eta_{1}(1855)$ may also be interpreted as the hybrid states $|\bar{q}qg;1^{-}1^{-+}\rangle$ and $|\bar{q}qg;0^{+}1^{-+}\rangle$, respectively. To differentiate these two assignments and to verify whether they are hybrid states or not, we propose to examine the $a_{1}(1260)\pi$ decay channel in future experiments.
hybrid state, exotic hadron, QCD sum rules
I Introduction
A single-gluon hybrid state is composed of one valence quark and one valence antiquark as well as one valence gluon. Especially, the hybrid states with $J^{PC}=0^{\pm-}/1^{-+}/2^{+-}/3^{-+}/\cdots$ are of particular interests, since these exotic quantum numbers can not be accessed by the conventional $\bar{q}q$ meson pdg ; Klempt:2007cp ; Amsler:2004ps ; Bugg:2004xu ; Meyer:2010ku ; Meyer:2015eta ; Chen:2016qju ; Briceno:2017max ; Ketzer:2019wmd ; Jin:2021vct ; Chen:2022asf . Experimental confirmation of the hybrid state is a direct test of QCD in the low energy sector. Recently, the BESIII collaboration performed a partial wave analysis of the $J/\psi\to\gamma\eta\eta^{\prime}$ decay, and observed the $\eta_{1}(1855)$ in the $\eta\eta^{\prime}$ mass spectrum with a statistical significance larger than $19\sigma$ BESIII:2022riz ; BESIII:2022iwi . This state has the exotic quantum number $I^{G}J^{PC}=0^{+}1^{-+}$. Its mass and width were measured to be
$\eta_{1}(1855)$ ::: $M=1855\pm 9^{+6}_{-1}{\rm~{}MeV}/c^{2}\,,$
$\Gamma=188\pm 18^{+3}_{-8}{\rm~{}MeV}\,.$
Besides the isoscalar state $\eta_{1}(1855)$, there are three isovector states, $\pi_{1}(1400)$ IHEP-Brussels-LosAlamos-AnnecyLAPP:1988iqi , $\pi_{1}(1600)$ E852:1998mbq , and $\pi_{1}(2015)$ E852:2004gpn , which have the exotic quantum number $I^{G}J^{PC}=1^{-}1^{-+}$. According to PDG, their masses and widths are pdg :
$\pi_{1}(1400)$ ::: $M=1354\pm 25{\rm~{}MeV}\,,$
$\Gamma=330\pm 35{\rm~{}MeV}\,;$
$\pi_{1}(1600)$ ::: $M=1661^{+15}_{-11}{\rm~{}MeV}\,,$
$\Gamma=240\pm 50{\rm~{}MeV}\,;$
$\pi_{1}(2015)$ ::: $M=2014\pm 20\pm 16{\rm~{}MeV}\,,$
$\Gamma=230\pm 32\pm 73{\rm~{}MeV}\,.$
The $\pi_{1}(1400)$ was observed in the $\eta\pi$ decay channel by several collaborations IHEP-Brussels-LosAlamos-AnnecyLAPP:1988iqi ; Aoyagi:1993kn ; E852:1997gvf ; VES:2001rwn ; CrystalBarrel:1998cfz ; CrystalBarrel:2019zqh . It was also observed in the $\rho\pi$ decay channel by OBELIX OBELIX:2004oio , but this was not confirmed by COMPASS COMPASS:2018uzl . The $\pi_{1}(1600)$ was observed in the $\rho\pi$, $\eta^{\prime}\pi$, $b_{1}(1235)\pi$, and $f_{1}(1285)\pi$ decay channels by several collaborations E852:1998mbq ; Khokhlov:2000tk ; Baker:2003jh ; COMPASS:2009xrl ; CLEO:2011upl , while the $\pi_{1}(2015)$ was observed in the $b_{1}(1235)\pi$ and $f_{1}(1285)\pi$ decay channels only in the BNL E852 experiments E852:2004gpn ; E852:2004rfa . In recent years the COMPASS and JPAC collaborations further examined the $\eta\pi$ and $\eta^{\prime}\pi$ decay channels COMPASS:2014vkj ; COMPASS:2018uzl ; JPAC:2018zyd ; COMPASS:2021ogp , and their results suggest that there is only one exotic $\pi_{1}$ resonance coupling to both the $\eta\pi$ and $\eta^{\prime}\pi$ channels, while there is no evidence for a second exotic resonance. Its mass and width were determined to be $1564\pm 24\pm 86$ MeV and $492\pm 54\pm 102$ MeV, respectively COMPASS:2014vkj .
In the past fifty years there have been a lot of theoretical investigations on the hybrid states, such as the MIT bag model Barnes:1977hg ; Hasenfratz:1980jv ; Chanowitz:1982qj , flux-tube model Isgur:1983wj ; Close:1994hc ; Page:1998gz ; Qiu:2022ktc , AdS/QCD model Andreev:2012hw ; Bellantuono:2014lra , lattice QCD Michael:1985ne ; Juge:2002br ; Lacock:1996ny ; MILC:1997usn ; Dudek:2013yja , QCD sum rules Reinders:1982hd ; Balitsky:1982ps ; Govaerts:1983ka ; Kisslinger:1995yw ; Jin:2002rw ; Li:2021fwk , and constituent gluon model Horn:1977rq ; Szczepaniak:2001rg ; Guo:2007sm , etc. However, their nature remains elusive since we still poorly understand the gluon degree of freedom. Experimentally, it is not easy to identify the hybrid states unambiguously, so there is currently no definite experimental evidence on their existence. Theoretically, it is also not easy to define the gluon degree of freedom, and a precise definition of the constituent gluon is still lacking. We refer to Refs. Horn:1977rq ; Szczepaniak:2001rg ; Guo:2007sm ; Coyne:1980zd ; Chanowitz:1980gu ; Barnes:1981ac ; Cornwall:1982zn ; Cho:2015rsa for discussions on how to construct glueballs and hybrid states using the constituent gluons.
There have been many theoretical calculations on the masses of the $J^{PC}=1^{-+}$ hybrid states Ebert:2009ub ; Kim:2008qh ; Ping:2009zza ; Kitazoe:1983xx ; Dudek:2010wm ; Dudek:2009qf ; Kisslinger:2009pw . For examples, their masses extracted from the lattice QCD simulations are about 1.7 GeV Hedditch:2005zf , 1.8 GeV Bernard:2003jd , and 2.0 GeV McNeile:1998cp , while their masses extracted from the flux tube model and the constituent gluon model are about 1.9 GeV Isgur:1984bm ; Burns:2006wz ; Iddir:2007dq . The QCD sum rule method has been widely applied to study the $J^{PC}=1^{-+}$ hybrid states in Refs. Govaerts:1984bk ; Chetyrkin:2000tj ; Huang:1998zj ; Latorre:1985tg ; Yang:2007cc ; Reinders:1981ww ; Narison:2009vj , and this method has also been applied to study the $J^{PC}=0^{+-}$ and $2^{+-}$ hybrid states in Refs. Ho:2018cat ; Wang:2023whb . We refer to Refs. Page:1996rj ; Isgur:1985vy ; DeViron:1984svx ; Frere:1988ac ; Zhu:1998sv ; Zhu:1999wg ; Godfrey:2002rp ; McNeile:2006bz ; Zhang:2002id ; Close:1994pr ; Afanasev:1997fp ; Szczepaniak:2001qz ; Shastry:2022mhk ; Chen:2022isv ; Yu:2022lwl ; Chen:2008ne ; Chen:2008qw ; Chen:2015moa ; Zhang:2019ykd ; Tang:2021zti ; Dong:2022cuw ; Yang:2022rck ; Wan:2022xkx ; Wang:2022sib ; Ji:2022blw ; Frere:2024wsf ; Barsbay:2024vjt for more theoretical studies. We also refer to our recent review Chen:2022asf as well as the reviews Klempt:2007cp ; Amsler:2004ps ; Bugg:2004xu ; Meyer:2010ku ; Meyer:2015eta ; Chen:2016qju ; Briceno:2017max ; Ketzer:2019wmd ; Jin:2021vct for detailed discussions.
In this paper we shall systematically investigate the single-gluon hybrid states through the QCD sum rule method. Especially, we shall concentrate on the hybrid states with the exotic quantum number $J^{PC}=1^{-+}$ and update the previous calculations on their mass spectrum Latorre:1985tg , given that some QCD parameters have been significantly changed in recent years. We shall also update our previous calculations on their decay properties Chen:2010ic ; Huang:2010dc ; Chen:2022qpd . Assuming their quark-gluon contents to be either $\bar{q}qg$ ($q=u/d$) or $\bar{s}sg$ and their isospin to be either $I=1$ or $I=0$, we shall calculate their masses and widths to be:
$M_{|\bar{q}qg;1^{-}1^{-+}\rangle}$ $=$ $1.67^{+0.15}_{-0.17}{\rm~{}GeV}\,,$
$\Gamma_{|\bar{q}qg;1^{-}1^{-+}\rangle}$ $=$ $530^{+540}_{-330}{\rm~{}MeV}\,,$
$M_{|\bar{q}qg;0^{+}1^{-+}\rangle}$ $=$ $1.67^{+0.15}_{-0.17}{\rm~{}GeV}\,,$
$\Gamma_{|\bar{q}qg;0^{+}1^{-+}\rangle}$ $=$ $120^{+160}_{-110}{\rm~{}MeV}\,,$
$M_{|\bar{s}sg;0^{+}1^{-+}\rangle}$ $=$ $1.84^{+0.14}_{-0.15}{\rm~{}GeV}\,,$
$\Gamma_{|\bar{s}sg;0^{+}1^{-+}\rangle}$ $=$ $100^{+110}_{-~{}80}{\rm~{}MeV}\,.$
This paper is organized as follows. In Sec. II we construct twenty-four single-gluon hybrid currents with various (exotic) quantum numbers, eighteen of which are used to perform QCD sum rule analyses in Sec. III. After performing numerical analyses in Sec. IV, we calculate the masses of forty-four single-gluon hybrid states with the quark-gluon contents $\bar{q}qg$ ($q=u/d$) and $\bar{s}sg$. The obtained results are summarized and discussed in Sec. V. Based on these results, we update our previous calculations on the decay properties of the $J^{PC}=1^{-+}$ hybrid states Chen:2010ic ; Huang:2010dc ; Chen:2022qpd , and the obtained results are also summarized in Sec. V.
II Single-gluon hybrid currents
In this section we systematically construct the single-gluon hybrid currents using the light quark field $q_{a}(x)$ and its dual field $\bar{q}_{a}(x)$ as well as the gluon field strength tensor $G^{n}_{\mu\nu}(x)$ and its dual field $\widetilde{G}^{n}_{\mu\nu}=G^{n,\rho\sigma}\times\epsilon_{\mu\nu\rho\sigma}/2$, with $a=1\cdots 3$ and $n=1\cdots 8$ the color indices, and $\mu\cdots\sigma$ the Lorentz indices. Generally speaking, we can construct the single-gluon hybrid currents by combining the color-octet quark-antiquark fields
$\bar{q}_{a}\lambda_{n}^{ab}\gamma_{5}q_{b}\,,\,\bar{q}_{a}\lambda_{n}^{ab}q_{b}\,,\,$
$\bar{q}_{a}\lambda_{n}^{ab}\gamma_{\mu}q_{b}\,,\,\bar{q}_{a}\lambda_{n}^{ab}\gamma_{\mu}\gamma_{5}q_{b}\,,\,$ (5)
$\bar{q}_{a}\lambda_{n}^{ab}\sigma_{\mu\nu}q_{b}\,,$
and the color-octet gluon fields
$G_{n}^{\alpha\beta}\,,\,\widetilde{G}_{n}^{\alpha\beta}\,,$ (6)
together with some Lorentz coefficients $\Gamma^{\mu\nu\cdots\alpha\beta}$.
Figure 1: Categorization of the single-gluon hybrid currents.
As summarized in Fig. 1, there are altogether twenty-four single-gluon hybrid currents, denoted as $J^{\cdots}_{J^{PC}}/\widetilde{J}^{\cdots}_{J^{PC}}$ with $J$ the total spin:
$J^{\alpha\beta}_{1^{--}}$ $=$ $\bar{q}_{a}\lambda_{n}^{ab}\gamma_{5}q_{b}~{}g_{s}G_{n}^{\alpha\beta}\,,$ (7)
$\widetilde{J}^{\alpha\beta}_{1^{+-}}$ $=$ $\bar{q}_{a}\lambda_{n}^{ab}\gamma_{5}q_{b}~{}g_{s}\widetilde{G}_{n}^{\alpha\beta}\,,$ (8)
$J^{\alpha\beta}_{1^{+-}}$ $=$ $\bar{q}_{a}\lambda_{n}^{ab}q_{b}~{}g_{s}G_{n}^{\alpha\beta}\,,$ (9)
$\widetilde{J}^{\alpha\beta}_{1^{--}}$ $=$ $\bar{q}_{a}\lambda_{n}^{ab}q_{b}~{}g_{s}\widetilde{G}_{n}^{\alpha\beta}\,,$ (10)
$J^{\mu}_{1^{-+}}$ $=$ $\bar{q}_{a}\lambda_{n}^{ab}\gamma_{\beta}q_{b}~{}g_{s}G_{n}^{\mu\beta}\,,$ (11)
$\widetilde{J}^{\mu}_{1^{++}}$ $=$ $\bar{q}_{a}\lambda_{n}^{ab}\gamma_{\beta}q_{b}~{}g_{s}\widetilde{G}_{n}^{\mu\beta}\,,$ (12)
$J^{\mu}_{1^{+-}}$ $=$ $\bar{q}_{a}\lambda_{n}^{ab}\gamma_{\beta}\gamma_{5}q_{b}~{}g_{s}G_{n}^{\mu\beta}\,,$ (13)
$\widetilde{J}^{\mu}_{1^{--}}$ $=$ $\bar{q}_{a}\lambda_{n}^{ab}\gamma_{\beta}\gamma_{5}q_{b}~{}g_{s}\widetilde{G}_{n}^{\mu\beta}\,,$ (14)
$J^{\mu,\alpha\beta}_{2^{-+}}$ $=$ $\bar{q}_{a}\lambda_{n}^{ab}\gamma^{\mu}q_{b}~{}g_{s}G_{n}^{\alpha\beta}\,,$ (15)
$\widetilde{J}^{\mu,\alpha\beta}_{2^{++}}$ $=$ $\bar{q}_{a}\lambda_{n}^{ab}\gamma^{\mu}q_{b}~{}g_{s}\widetilde{G}_{n}^{\alpha\beta}\,,$ (16)
$J^{\mu,\alpha\beta}_{2^{+-}}$ $=$ $\bar{q}_{a}\lambda_{n}^{ab}\gamma^{\mu}\gamma_{5}q_{b}~{}g_{s}G_{n}^{\alpha\beta}\,,$ (17)
$\widetilde{J}^{\mu,\alpha\beta}_{2^{--}}$ $=$ $\bar{q}_{a}\lambda_{n}^{ab}\gamma^{\mu}\gamma_{5}q_{b}~{}g_{s}\widetilde{G}_{n}^{\alpha\beta}\,,$ (18)
$J_{0^{++}}$ $=$ $\bar{q}_{a}\lambda_{n}^{ab}\sigma_{\mu\nu}q_{b}~{}g_{s}G_{n}^{\mu\nu}\,,$ (19)
$\widetilde{J}_{0^{-+}}$ $=$ $\bar{q}_{a}\lambda_{n}^{ab}\sigma_{\mu\nu}q_{b}~{}g_{s}\widetilde{G}_{n}^{\mu\nu}\,,$ (20)
$J_{0^{-+}}$ $=$ $\bar{q}_{a}\lambda_{n}^{ab}\sigma_{\mu\nu}\gamma_{5}q_{b}~{}g_{s}G_{n}^{\mu\nu}\,,$ (21)
$\widetilde{J}_{0^{++}}$ $=$ $\bar{q}_{a}\lambda_{n}^{ab}\sigma_{\mu\nu}\gamma_{5}q_{b}~{}g_{s}\widetilde{G}_{n}^{\mu\nu}\,,$ (22)
$J^{\alpha\beta}_{1^{++}}$ $=$ $\mathcal{A}[\bar{q}_{a}\lambda_{n}^{ab}\sigma^{\alpha\mu}q_{b}~{}g_{s}G_{n,\mu}^{\beta}]\,,$ (23)
$\widetilde{J}^{\alpha\beta}_{1^{-+}}$ $=$ $\mathcal{A}[\bar{q}_{a}\lambda_{n}^{ab}\sigma^{\alpha\mu}q_{b}~{}g_{s}\widetilde{G}_{n,\mu}^{\beta}]\,,$ (24)
$J^{\alpha\beta}_{1^{-+}}$ $=$ $\mathcal{A}[\bar{q}_{a}\lambda_{n}^{ab}\sigma^{\alpha\mu}\gamma_{5}q_{b}~{}g_{s}G_{n,\mu}^{\beta}]\,,$ (25)
$\widetilde{J}^{\alpha\beta}_{1^{++}}$ $=$ $\mathcal{A}[\bar{q}_{a}\lambda_{n}^{ab}\sigma^{\alpha\mu}\gamma_{5}q_{b}~{}g_{s}\widetilde{G}_{n,\mu}^{\beta}]\,,$ (26)
$J^{\alpha_{1}\beta_{1},\alpha_{2}\beta_{2}}_{2^{++}}$ $=$ $\mathcal{S}[\bar{q}_{a}\lambda_{n}^{ab}\sigma^{\alpha_{1}\beta_{1}}q_{b}~{}g_{s}G_{n}^{\alpha_{2}\beta_{2}}]\,,$ (27)
$\widetilde{J}^{\alpha_{1}\beta_{1},\alpha_{2}\beta_{2}}_{2^{-+}}$ $=$ $\mathcal{S}[\bar{q}_{a}\lambda_{n}^{ab}\sigma^{\alpha_{1}\beta_{1}}q_{b}~{}g_{s}\widetilde{G}_{n}^{\alpha_{2}\beta_{2}}]\,,$ (28)
$J^{\alpha_{1}\beta_{1},\alpha_{2}\beta_{2}}_{2^{-+}}$ $=$ $\mathcal{S}[\bar{q}_{a}\lambda_{n}^{ab}\sigma^{\alpha_{1}\beta_{1}}\gamma_{5}q_{b}~{}g_{s}G_{n}^{\alpha_{2}\beta_{2}}]\,,$ (29)
$\widetilde{J}^{\alpha_{1}\beta_{1},\alpha_{2}\beta_{2}}_{2^{++}}$ $=$ $\mathcal{S}[\bar{q}_{a}\lambda_{n}^{ab}\sigma^{\alpha_{1}\beta_{1}}\gamma_{5}q_{b}~{}g_{s}\widetilde{G}_{n}^{\alpha_{2}\beta_{2}}]\,.$ (30)
Especially, we shall concentrate on the fifth current $J^{\mu}_{1^{-+}}$ with the exotic quantum number $J^{PC}=1^{-+}$. In the above expressions $\{\alpha\beta\}/\{\alpha_{1}\beta_{1}\}/\{\alpha_{2}\beta_{2}\}$ are antisymmetric Lorentz pairs. The four currents $J^{\mu,\alpha\beta}_{2^{-+}}$, $\widetilde{J}^{\mu,\alpha\beta}_{2^{++}}$, $J^{\mu,\alpha\beta}_{2^{+-}}$, and $\widetilde{J}^{\mu,\alpha\beta}_{2^{--}}$ all contain three Lorentz indices with the mixed symmetry, so their spin-2 components can not be easily extracted. We shall not investigate these four currents in the present study, and we refer to Ref. Wang:2023whb for detailed discussions. The symbol $\mathcal{A}[\cdots]$ represents anti-symmetrization in the set $\{\alpha\beta\}$, which can be done by multiplying the projection operator
$\Gamma_{\alpha^{\prime}\beta^{\prime};\alpha\beta}$ $=$ $g_{\alpha^{\prime}\alpha}g_{\beta^{\prime}\beta}-g_{\beta^{\prime}\alpha}g_{\alpha^{\prime}\beta}\,.$ (31)
The symbol $\mathcal{S}[\cdots]$ represents symmetrization and subtracting trace terms in the two sets $\{\alpha_{1}\alpha_{2}\}$ and $\{\beta_{1}\beta_{2}\}$ simultaneously. In the present study we only need to investigate its leading spin-2 component, which can be done by multiplying the projection operator
$\Gamma^{\prime}_{\alpha_{1}^{\prime}\beta_{1}^{\prime},\alpha_{2}^{\prime}\beta_{2}^{\prime};\alpha_{1}\beta_{1},\alpha_{2}\beta_{2}}$
$=$ $(g_{\alpha_{1}^{\prime}\alpha_{1}}g_{\beta_{1}^{\prime}\beta_{1}}-g_{\beta_{1}^{\prime}\alpha_{1}}g_{\alpha_{1}^{\prime}\beta_{1}})(g_{\alpha_{2}^{\prime}\alpha_{2}}g_{\beta_{2}^{\prime}\beta_{2}}-g_{\beta_{2}^{\prime}\alpha_{2}}g_{\alpha_{2}^{\prime}\beta_{2}})$
$+$ $(g_{\alpha_{1}^{\prime}\alpha_{2}}g_{\beta_{1}^{\prime}\beta_{1}}-g_{\beta_{1}^{\prime}\alpha_{2}}g_{\alpha_{1}^{\prime}\beta_{1}})(g_{\alpha_{2}^{\prime}\alpha_{1}}g_{\beta_{2}^{\prime}\beta_{2}}-g_{\beta_{2}^{\prime}\alpha_{1}}g_{\alpha_{2}^{\prime}\beta_{2}})$
$+$ $(g_{\alpha_{1}^{\prime}\alpha_{1}}g_{\beta_{1}^{\prime}\beta_{2}}-g_{\beta_{1}^{\prime}\alpha_{1}}g_{\alpha_{1}^{\prime}\beta_{2}})(g_{\alpha_{2}^{\prime}\alpha_{2}}g_{\beta_{2}^{\prime}\beta_{1}}-g_{\beta_{2}^{\prime}\alpha_{2}}g_{\alpha_{2}^{\prime}\beta_{1}})$
$+$ $(g_{\alpha_{1}^{\prime}\alpha_{2}}g_{\beta_{1}^{\prime}\beta_{2}}-g_{\beta_{1}^{\prime}\alpha_{2}}g_{\alpha_{1}^{\prime}\beta_{2}})(g_{\alpha_{2}^{\prime}\alpha_{1}}g_{\beta_{2}^{\prime}\beta_{1}}-g_{\beta_{2}^{\prime}\alpha_{1}}g_{\alpha_{2}^{\prime}\beta_{1}})$
$+$ $\cdots$.⋯ .
where $\cdots$ contains the irrelevant terms.
Before performing QCD sum rule analyses, we separately discuss the Lorentz structures of the above single-gluon hybrid currents as follows:
  • Due to the formulae $\sigma_{\mu\nu}\gamma_{5}=\epsilon_{\mu\nu\rho\sigma}\sigma^{\rho\sigma}\times i/2$, the two currents $J_{0^{++}}$ and $\widetilde{J}_{0^{++}}$ are equivalent, and the other two currents $J_{0^{-+}}$ and $\widetilde{J}_{0^{-+}}$ are also equivalent. Hence, we shall only study the currents $J_{0^{++}}$ and $J_{0^{-+}}$.
  • The current $J^{\alpha\beta}_{1^{--}}$, with two antisymmetric Lorentz indices $\{\alpha\beta\}$, contains both the $J^{PC}=1^{--}$ and $1^{+-}$ components, so it couples to both the $J^{PC}=1^{--}$ and $1^{+-}$ states through
    $\langle 0|J^{\alpha\beta}_{1^{--}}|X_{1^{--}}\rangle$ $=$ $if_{1^{--}}\epsilon^{\alpha\beta\mu\nu}\epsilon_{\mu}q_{\nu}\,,$ (33)
    $\langle 0|J^{\alpha\beta}_{1^{--}}|\widetilde{X}_{1^{+-}}\rangle$ $=$ $i\widetilde{f}_{1^{+-}}(q^{\alpha}\epsilon^{\beta}-q^{\beta}\epsilon^{\alpha})\,,$ (34)
    where $f_{1^{--}}$ and $\widetilde{f}_{1^{+-}}$ are two decay constants. Given the Lorentz structures of Eq. (33) and Eq. (34) to be totally different, we can clearly separate the two states $X_{1^{--}}$ and $\widetilde{X}_{1^{+-}}$ at the hadron level, i.e., we can isolate $X_{1^{--}}$ by investigating the two-point correlation function containing
    $\langle 0|J^{\alpha\beta}_{1^{--}}|X_{1^{--}}\rangle\langle X_{1^{--}}|J^{\alpha^{\prime}\beta^{\prime}\dagger}_{1^{--}}|0\rangle$
    $=$ $f_{1^{--}}^{2}~{}\epsilon^{\alpha\beta\mu\nu}\epsilon_{\mu}q_{\nu}~{}\epsilon^{\alpha^{\prime}\beta^{\prime}\mu^{\prime}\nu^{\prime}}\epsilon^{*}_{\mu^{\prime}}q_{\nu^{\prime}}$
    $=$ $-f_{1^{--}}^{2}~{}q^{2}~{}\left(g^{\alpha\alpha^{\prime}}g^{\beta\beta^{\prime}}-g^{\alpha\beta^{\prime}}g^{\beta\alpha^{\prime}}\right)+\cdots\,,$
    while the correlation function of $\widetilde{X}_{1^{+-}}$ dose not contain the above coefficient. It is not so easy to isolate $\widetilde{X}_{1^{+-}}$ from $J^{\alpha\beta}_{1^{--}}$, but instead, we can study the dual current $\widetilde{J}^{\alpha\beta}_{1^{+-}}$ that couples to $X_{1^{--}}$ and $\widetilde{X}_{1^{+-}}$ in the opposite ways. According to the above analysis, we shall study the single-gluon hybrid currents $J^{\alpha\beta}_{1^{--}}/\widetilde{J}^{\alpha\beta}_{1^{+-}}/J^{\alpha\beta}_{1^{+-}}/\widetilde{J}^{\alpha\beta}_{1^{--}}/J^{\alpha\beta}_{1^{++}}/\widetilde{J}^{\alpha\beta}_{1^{-+}}/J^{\alpha\beta}_{1^{-+}}/\widetilde{J}^{\alpha\beta}_{1^{++}}$ to investigate the single-gluon hybrid states of $J^{PC}=1^{--}/1^{+-}/1^{+-}/1^{--}/1^{++}/1^{-+}/1^{-+}/1^{++}$, respectively.
  • The four currents $J^{\alpha_{1}\beta_{1},\alpha_{2}\beta_{2}}_{2^{++}}$, $\widetilde{J}^{\alpha_{1}\beta_{1},\alpha_{2}\beta_{2}}_{2^{-+}}$, $J^{\alpha_{1}\beta_{1},\alpha_{2}\beta_{2}}_{2^{-+}}$, and $\widetilde{J}^{\alpha_{1}\beta_{1},\alpha_{2}\beta_{2}}_{2^{++}}$ all contain various $J^{PC}$ components and so couple to many states. We shall only investigate the four $J=2$ states of $J^{PC}=2^{++}/2^{-+}/2^{-+}/2^{++}$ by calculating the highest-order correlation functions, e.g.:
    $\Pi_{2^{++}}^{\alpha_{1}\beta_{1},\alpha_{2}\beta_{2};\alpha_{1}^{\prime}\beta_{1}^{\prime},\alpha_{2}^{\prime}\beta_{2}^{\prime}}(q^{2})$
    $=$ $i\int d^{4}xe^{iqx}\langle 0|{\bf T}[J^{\alpha_{1}\beta_{1},\alpha_{2}\beta_{2}}_{2^{++}}(x)J^{\alpha_{1}^{\prime}\beta_{1}^{\prime},\alpha_{2}^{\prime}\beta_{2}^{\prime}\dagger}_{2^{++}}(0)]|0\rangle$
    $=$ $\mathcal{S}^{\prime}[g^{\alpha_{1}\alpha_{1}^{\prime}}g^{\beta_{1}\beta_{1}^{\prime}}g^{\alpha_{2}\alpha_{2}^{\prime}}g^{\beta_{2}\beta_{2}^{\prime}}]~{}\Pi_{2^{++}}(q^{2})+\cdots\,,$
    where $\mathcal{S}^{\prime}[\cdots]$ represents symmetrization and subtracting trace terms in the four sets $\{\alpha_{1}\alpha_{2}\}$, $\{\beta_{1}\beta_{2}\}$, $\{\alpha_{1}^{\prime}\alpha_{2}^{\prime}\}$, and $\{\beta_{1}^{\prime}\beta_{2}^{\prime}\}$ simultaneously. The correlation function $\Pi_{2^{++}}(q^{2})$ is contributed only by the $J^{PC}=2^{++}$ component, while $\cdots$ contains the contributions from all the $J^{PC}$ components coupling to $J^{\alpha_{1}\beta_{1},\alpha_{2}\beta_{2}}_{2^{++}}$. However, we do not know the explicit expression of the rearrangement $\mathcal{S}^{\prime}[\cdots]$, so we are not able to calculate the decay constants of these four currents.
  • The current $J^{\mu}_{1^{-+}}$ contains both the $J^{PC}=1^{-+}$ and $0^{++}$ components, so it couples to both the $J^{PC}=1^{+-}$ and $0^{++}$ states through
    $\langle 0|J^{\mu}_{1^{-+}}|X_{1^{-+}}\rangle$ $=$ $\epsilon^{\mu}f_{1^{-+}}\,,$ (37)
    $\langle 0|J^{\mu}_{1^{-+}}|X_{0^{++}}\rangle$ $=$ $q^{\mu}f_{0^{++}}\,,$ (38)
    where $f_{1^{-+}}$ and $f_{0^{++}}$ are two decay constants. We can clearly separate the two states $X_{1^{-+}}$ and $X_{0^{++}}$ at the hadron level by calculating both $\Pi_{1^{-+}}(q^{2})$ and $\Pi_{0^{++}}(q^{2})$ of the two-point correlation function
    $\Pi_{1^{-+}}^{\mu\nu}(q^{2})$
    $\equiv$ $i\int d^{4}xe^{iqx}\langle 0|{\bf T}[J^{\mu}_{1^{-+}}(x)J^{\nu\dagger}_{1^{-+}}(0)]|0\rangle$
    $=$ $(g^{\mu\nu}-q^{\mu}q^{\nu}/q^{2})~{}\Pi_{1^{-+}}(q^{2})+(q^{\mu}q^{\nu}/q^{2})~{}\Pi_{0^{++}}(q^{2})\,.$
    Similarly, we shall study the single-gluon hybrid currents $\widetilde{J}^{\mu}_{1^{++}}/J^{\mu}_{1^{+-}}/\widetilde{J}^{\mu}_{1^{--}}$ to investigate the single-gluon hybrid states of both $J^{PC}=1^{++}/1^{+-}/1^{--}$ and $J^{PC}=0^{-+}/0^{--}/0^{+-}$.
III QCD sum rule analyses
In this section we use the single-gluon hybrid currents given in Eqs. (7-14) and Eqs. (19-30) to perform QCD sum rule analyses. We use the current $J^{\mu}_{1^{-+}}$ given in Eq. (11) as an example. Based on Eq. (37) and Eq. (II), we study its two-point correlation function
$\Pi_{1^{-+}}^{\mu\nu}(q^{2})$ $=$ $i\int d^{4}xe^{iqx}\langle 0|{\bf T}[J^{\mu}_{1^{-+}}(x)J^{\nu\dagger}_{1^{-+}}(0)]|0\rangle$ (40)
$=$ $(g^{\mu\nu}-q^{\mu}q^{\nu}/q^{2})~{}\Pi_{1^{-+}}(q^{2})+\cdots,$
at both the hadron and quark-gluon levels, where $\Pi_{1^{-+}}(q^{2})$ is contributed by the $J^{PC}=1^{-+}$ state $X_{1^{-+}}$, and $\cdots$ is contributed by the $J^{PC}=0^{++}$ state $X_{0^{++}}$.
We use the dispersion relation to express $\Pi_{1^{-+}}(q^{2})$ as
[…]
where $\rho_{1^{-+}}(s)\equiv{\rm Im}\Pi_{1^{-+}}(s)/\pi$ is the spectral density, and $s_{<}=4m_{q}^{2}$ is the physical threshold.
At the hadron level, we parameterize $\rho_{1^{-+}}^{\rm phen}(s)$ as one pole dominance for the state $X_{1^{-+}}$ together with a continuum contribution:
$\rho_{1^{-+}}^{\rm phen}(s)\times(g^{\mu\nu}-q^{\mu}q^{\nu}/q^{2})$
$\equiv$ $\sum_{n}\delta(s-M^{2}_{n})\langle 0|J^{\mu}_{1^{-+}}|n\rangle\langle n|J^{\nu\dagger}_{1^{-+}}|0\rangle$
$=$ $f_{1^{-+}}^{2}\delta(s-M^{2}_{1^{-+}})\times(g^{\mu\nu}-q^{\mu}q^{\nu}/q^{2})+\rm{continuum}\,.$
At the quark-gluon level, we calculate $\rho_{1^{-+}}^{\rm OPE}(s)$ through the method of operator product expansion (OPE). After performing the Borel transformation at both the hadron and quark-gluon levels, we use $\rho_{1^{-+}}^{\rm OPE}(s)$ above the threshold value $s_{0}$ to approximate the continuum, and derive the sum rule equation
$\Pi_{1^{-+}}(s_{0},M_{B}^{2})$$\equiv$ $f^{2}_{1^{-+}}e^{-M_{1^{-+}}^{2}/M_{B}^{2}}$
[…]
which can be used to further derive
[…]
$f^{2}_{1^{-+}}(s_{0},M_{B})$ $=$ $\Pi_{1^{-+}}(s_{0},M_{B}^{2})\times e^{M_{1^{-+}}^{2}/M_{B}^{2}}.$ (45)
(a) (a)
(b) (b–1)
(c) (b–2)
(d) (b–3)
(e) (c–1)
(f) (c–2)
(g) (d–1)
(h) (d–2)
(i) (d–3)
(j) (d–4)
(k) (e–1)
(l) (e–2)
Figure 2: Feynman diagrams for the single-gluon hybrid state: (a) and (b–i) are proportional to $\alpha_{s}\times g_{s}^{0}$; (c–i) and (d–i) are proportional to $\alpha_{s}\times g_{s}^{1}$; (e–i) are proportional to $\alpha_{s}\times g_{s}^{2}$.
In the present study we have considered the Feynman diagrams depicted in Fig. 2, and calculated $\rho_{\rm OPE}(s)$ up to the dimension eight condensates. The gluon field strength tensor $G^{n}_{\mu\nu}$ is defined as
$G^{n}_{\mu\nu}=\partial_{\mu}A_{\nu}^{n}-\partial_{\nu}A_{\mu}^{n}+g_{s}f^{npq}A_{p,\mu}A_{q,\nu}\,,$ (46)
which can be separated into the former two terms (represented by the single-gluon-line) and the third term (represented by the double-gluon-line with a red vertex, e.g., see the diagram depicted in Fig. 2(c–3)). In the calculations we have considered the perturbative term, the quark condensates, the quark-gluon mixed condensates, the two-gluon condensate, the three-gluon condensate, and their combinations. We have taken into account all the diagrams proportional to $\alpha_{s}\times g_{s}^{0}$ and $\alpha_{s}\times g_{s}^{1}$, but we have taken into account only three diagrams proportional to $\alpha_{s}\times g_{s}^{2}$. We summarize the obtained OPE spectral densities in Appendix A. Especially, the one extracted from the current $J^{\mu}_{1^{-+}}$ with the quark-gluon content $\bar{s}sg$ is
$\Pi^{\mu}_{1^{-+}}{\left({M_{B}^{2},{s_{0}}}\right)}$
$=$ $\int_{4m_{s}^{2}}^{{s_{0}}}{(\frac{{{s^{3}}{\alpha_{s}}}}{{60{\pi^{3}}}}}-\frac{{m_{s}^{2}{s^{2}}{\alpha_{s}}}}{{3{\pi^{3}}}}+s(\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle}}{{36{\pi^{2}}}}$
$+\frac{{13\left\langle{{\alpha_{s}}GG}\right\rangle{\alpha_{s}}}}{{432{\pi^{3}}}}+\frac{{8{m_{s}}\left\langle{\bar{s}s}\right\rangle{\alpha_{s}}}}{{9\pi}})+\frac{{\left\langle{g_{s}^{3}{G^{3}}}\right\rangle}}{{32{\pi^{2}}}}$
$-\frac{{3\left\langle{{\alpha_{s}}GG}\right\rangle m_{s}^{2}{\alpha_{s}}}}{{64{\pi^{3}}}}-\frac{{3{m_{s}}\left\langle{{g_{s}}\bar{s}\sigma Gs}\right\rangle{\alpha_{s}}}}{{4\pi}})\times{e^{-s/M_{B}^{2}}}ds$
$+$ $(\frac{{{{\left\langle{{\alpha_{s}}GG}\right\rangle}^{2}}}}{{3456{\pi^{2}}}}-\frac{{\left\langle{g_{s}^{3}{G^{3}}}\right\rangle m_{s}^{2}}}{{16{\pi^{2}}}}-\frac{2}{9}\left\langle{{\alpha_{s}}GG}\right\rangle{m_{s}}\left\langle{\bar{s}s}\right\rangle$
$+\frac{{11\pi\left\langle{\bar{s}s}\right\rangle\left\langle{{g_{s}}\bar{s}\sigma Gs}\right\rangle{\alpha_{s}}}}{9})\,.$
The one with the quark-gluon content $\bar{q}qg$ ($q=u/d$) can be easily derived by replacing $m_{s}\rightarrow 0$, $\left\langle{\bar{s}s}\right\rangle\rightarrow\left\langle{\bar{q}q}\right\rangle$, and $\left\langle{{g_{s}}\bar{s}\sigma Gs}\right\rangle\rightarrow\left\langle{{g_{s}}\bar{q}\sigma Gq}\right\rangle$. Note that we do not differentiate the up and down quarks in the calculations, so the states in the same isospin multiplet have the same extracted hadron mass, e.g., the two $\bar{q}qg$ states ($q=u/d$) with the quantum numbers $I^{G}J^{PC}=0^{+}1^{-+}$ and $1^{-}1^{-+}$, coupled by the same current $J^{\mu}_{1^{-+}}$, have the same extracted hadron mass.
IV Numerical Analyses
In this section we use the OPE spectral densities given in Eq. (III) as well as those listed in Appendix A to perform numerical analyses. We use the spectral density $\rho^{\bar{s}sg}_{1^{-+}}(s)$ given in Eq. (III) as an example. It is extracted from the current $J^{\mu}_{1^{-+}}$ with the quark-gluon content $\bar{s}sg$, which couples to the state
$X_{1^{-+}}\equiv|\bar{s}sg;1^{-+}\rangle\,.$ (48)
We shall use the following values for various QCD parameters at the renormalization scale 2 GeV and the QCD scale $\Lambda_{\rm QCD}=300$ MeV pdg ; Ovchinnikov:1988gk ; Yang:1993bp ; Ellis:1996xc ; Ioffe:2002be ; Jamin:2002ev ; Gimenez:2005nt ; Narison:2011xe ; Narison:2018dcr :
$\alpha_{s}(Q^{2})$ $=$ ${4\pi\over 11\ln(Q^{2}/\Lambda_{\rm QCD}^{2})}\,,$
$\langle\bar{q}q\rangle$ $=$ $-(0.240\pm 0.010)^{3}\mbox{ GeV}^{3}\,,$
$\langle\bar{s}s\rangle$ $=$ $(0.8\pm 0.1)\times\langle\bar{q}q\rangle\,,$
$\langle g_{s}\bar{q}\sigma Gq\rangle$ $=$ $(0.8\pm 0.2)\times\langle\bar{q}q\rangle\mbox{ GeV}^{2}\,,$ (49)
$\langle g_{s}\bar{s}\sigma Gs\rangle$ $=$ $(0.8\pm 0.2)\times\langle\bar{s}s\rangle\,,$
$\langle\alpha_{s}GG\rangle$ $=$ $(6.35\pm 0.35)\times 10^{-2}\mbox{ GeV}^{4}\,,$
$\langle g_{s}^{3}G^{3}\rangle$ $=$ $(8.2\pm 1.0)\times\langle\alpha_{s}GG\rangle\mbox{ GeV}^{2}\,,$
$m_{s}$ $=$ $93^{+11}_{-~{}5}\mbox{ MeV}\,.$
Note that the value of the gluon condensate $\langle\alpha_{s}GG\rangle$ is taken from Ref. Narison:2018dcr , which was written in 2018.
The mass $M_{1^{-+}}$ calculated by Eq. (44) depends on two free parameters: the Borel mass $M_{B}$ and the threshold value $s_{0}$. We shall determine their proper working regions through three criteria: a) the sufficiently good convergence of OPE, b) the sufficiently large pole contribution, and c) the sufficiently weak dependence of the mass $M_{1^{-+}}$ on these two parameters.
In order to have the sufficiently good convergence of OPE, we require the $\alpha_{s}\times g_{s}^{2}$ terms to be less than 5% and the $D=6+8$ terms to be less than 10%:
$\mbox{CVG}_{A}$$\equiv$ $\left|\frac{\Pi^{g_{s}^{n=4}}(\infty,M_{B}^{2})}{\Pi(\infty,M_{B}^{2})}\right|\leq 5\%\,,$ (50)
$\mbox{CVG}_{B}$$\equiv$ $\left|\frac{\Pi^{{\rm D=6+8}}(\infty,M_{B}^{2})}{\Pi(\infty,M_{B}^{2})}\right|\leq 10\%\,.$ (51)
As depicted in Fig. 3, we determine the minimum Borel mass to be $M_{B}^{2}\geq 2.26$ GeV2, when setting $s_{0}=6.2$ GeV2.
In order to have the sufficiently large pole contribution, we require
$\mbox{PC}\equiv\left|\frac{\Pi(s_{0},M_{B}^{2})}{\Pi(\infty,M_{B}^{2})}\right|\geq 40\%\,.$ (52)
As depicted in Fig. 3, we determine the maximum Borel mass to be $M_{B}^{2}\leq 2.54$ GeV2, when setting $s_{0}=6.2$ GeV2.
Figure 3: CVGA/B and PC with respect to the Borel mass $M_{B}$, when setting $s_{0}=6.2$ GeV2. These curves are obtained using the spectral density $\rho^{\bar{s}sg}_{1^{-+}}(s)$ extracted from the current $J^{\mu}_{1^{-+}}$ with the quark-gluon content $\bar{s}sg$.
Refer to captionRefer to caption
Figure 4: Mass of the single-gluon hybrid state $|\bar{s}sg;1^{-+}\rangle$ with respect to the threshold value $s_{0}$ (left) and the Borel mass $M_{B}$ (right). In the left panel the dotted, solid, and dashed curves are obtained by setting $M_{B}^{2}=2.26$ GeV2, $2.40$ GeV2, and $2.54$ GeV2, respectively. In the right panel the dotted, solid, and dashed curves are obtained by setting $s_{0}=5.2$ GeV2, $6.2$ GeV2, and $7.2$ GeV2, respectively. These curves are obtained using the spectral density $\rho^{\bar{s}sg}_{1^{-+}}(s)$ extracted from the current $J^{\mu}_{1^{-+}}$ with the quark-gluon content $\bar{s}sg$.
[…]
$M_{|\bar{s}sg;1^{-+}\rangle}$ $=$ $1.84^{+0.14}_{-0.15}{\rm~{}GeV}\,,$ (53)
$f_{|\bar{s}sg;1^{-+}\rangle}$ $=$ $0.300^{+0.063}_{-0.058}{\rm~{}GeV}\,.$ (54)
[…]
Similarly, we perform numerical analyses using the other single-gluon hybrid currents with the quark-gluon contents $\bar{q}qg$ ($q=u/d$) and $\bar{s}sg$. The obtained results are summarized in Table 1 and Table 2. Especially, we use the current $J^{\mu}_{1^{-+}}$ with the quark-gluon content $\bar{q}qg$ ($q=u/d$) to calculate the mass and decay constant of the single-gluon hybrid state $|\bar{q}qg;1^{-+}\rangle$ to be
$M_{|\bar{q}qg;1^{-+}\rangle}$ $=$ $1.67^{+0.15}_{-0.17}{\rm~{}GeV}\,,$ (55)
$f_{|\bar{q}qg;1^{-+}\rangle}$ $=$ $0.243^{+0.057}_{-0.052}{\rm~{}GeV}\,.$ (56)
V Summary and Discussions
Table 1: QCD sum rule results for the single-gluon hybrid states $|\bar{q}qg;J^{PC}\rangle$, extracted from the single-gluon hybrid currents given in Eqs. (7-14) and Eqs. (19-30) with the quark-gluon contents $\bar{q}qg$ ($q=u/d$). The results for the isoscalar state $|\bar{q}qg;0^{G}J^{PC}\rangle$ and the isovector state $|\bar{q}qg;1^{G}J^{PC}\rangle$ within the same isospin multiplet are the same as each other.
 State [$J^{PC}$]     Current   $s_{0}^{min}~{}[{\rm GeV}^{2}]$ Working Regions   Pole [%]   Mass [GeV]    Decay Constant
  $M_{B}^{2}~{}[{\rm GeV}^{2}]$   $s_{0}~{}[{\rm GeV}^{2}]$
$|\bar{q}qg;1^{--}\rangle$ $J^{\alpha\beta}_{1^{--}}$ 4.2 2.032.032.032.032.482.482.482.48 $5.5$ 4040404054545454 $1.80^{+0.13}_{-0.16}$ $0.051^{+0.004}_{-0.004}$ GeV3
$|\bar{q}qg;1^{+-}\rangle$ $\tilde{J}^{\alpha\beta}_{1^{+-}}$ 16.2 3.613.613.613.614.584.584.584.58 $18.0$ 4040404053535353 $4.05^{+0.24}_{-0.12}$ $0.063^{+0.020}_{-0.020}$ GeV3
$|\bar{q}qg;1^{+-}\rangle$ $J^{\alpha\beta}_{1^{+-}}$ 5.0 2.292.292.292.292.452.452.452.45 $5.5$ 4040404045454545 $1.84^{+0.12}_{-0.14}$ $0.049^{+0.004}_{-0.004}$ GeV3
$|\bar{q}qg;1^{--}\rangle$ $\tilde{J}^{\alpha\beta}_{1^{--}}$ 16.3 3.523.523.523.524.564.564.564.56 $18.0$ 4040404053535353 $4.09^{+0.29}_{-0.14}$ $0.064^{+0.021}_{-0.020}$ GeV3
$|\bar{q}qg;0^{++}\rangle$ $J^{\mu\rightarrow 0}_{1^{-+}}$ 20.6 5.115.115.115.116.596.596.596.59 $24.0$ 4040404056565656 $4.45^{+0.22}_{-0.17}$ $0.124^{+0.032}_{-0.036}$ GeV3
$|\bar{q}qg;0^{-+}\rangle$ $\tilde{J}^{\mu\rightarrow 0}_{1^{++}}$ 7.7 3.583.583.583.583.813.813.813.81 $8.5$ 4040404045454545 $2.14^{+0.17}_{-0.19}$ $0.105^{+0.005}_{-0.004}$ GeV3
$|\bar{q}qg;0^{--}\rangle$ $J^{\mu\rightarrow 0}_{1^{+-}}$ 21.6 5.485.485.485.486.526.526.526.52 $24.0$ 4040404050505050 $4.49^{+0.21}_{-0.14}$ $0.123^{+0.032}_{-0.037}$ GeV3
$|\bar{q}qg;0^{+-}\rangle$ $\tilde{J}^{\mu\rightarrow 0}_{1^{--}}$ 7.1 3.323.323.323.323.733.733.733.73 $8.5$ 4040404049494949 $2.16^{+0.16}_{-0.19}$ $0.100^{+0.005}_{-0.005}$ GeV3
$|\bar{q}qg;1^{-+}\rangle$ $J^{\mu}_{1^{-+}}$ 4.8 2.192.192.192.192.282.282.282.28 $5.2$ 4040404043434343 $1.67^{+0.15}_{-0.17}$ $0.243^{+0.057}_{-0.052}$ GeV4
$|\bar{q}qg;1^{++}\rangle$ $\tilde{J}^{\mu}_{1^{++}}$ 13.8 3.593.593.593.594.104.104.104.10 $15.0$ 4040404048484848 $3.54^{+0.16}_{-0.12}$ $1.370^{+0.494}_{-0.450}$ GeV4
$|\bar{q}qg;1^{+-}\rangle$ $J^{\mu}_{1^{+-}}$ 4.6 2.102.102.102.102.272.272.272.27 $5.2$ 4040404046464646 $1.68^{+0.14}_{-0.16}$ $0.242^{+0.055}_{-0.051}$ GeV4
$|\bar{q}qg;1^{--}\rangle$ $\tilde{J}^{\mu}_{1^{--}}$ 13.7 3.573.573.573.574.104.104.104.10 $15.0$ 4040404049494949 $3.53^{+0.16}_{-0.12}$ $1.366^{+0.493}_{-0.450}$ GeV4
$|\bar{q}qg;0^{++}\rangle$ $J_{0^{++}}$ 11.1 3.483.483.483.483.913.913.913.91 $12.5$ 4040404049494949 $2.94^{+0.20}_{-0.25}$ $2.893^{+1.029}_{-0.948}$ GeV4
$|\bar{q}qg;0^{-+}\rangle$ $J_{0^{-+}}$ 11.1 3.473.473.473.473.923.923.923.92 $12.5$ 4040404049494949 $2.93^{+0.20}_{-0.25}$ $2.882^{+1.026}_{-0.945}$ GeV4
$|\bar{q}qg;1^{++}\rangle$ $J^{\alpha\beta}_{1^{++}}$ 5.8 1.841.841.841.842.062.062.062.06 $6.5$ 4040404048484848 $2.11^{+0.17}_{-0.21}$ $0.056^{+0.012}_{-0.013}$ GeV3
$|\bar{q}qg;1^{-+}\rangle$ $\tilde{J}^{\alpha\beta}_{1^{-+}}$ 5.5 1.811.811.811.812.002.002.002.00 $6.2$ 4040404048484848 $2.00^{+0.13}_{-0.16}$ $0.055^{+0.007}_{-0.008}$ GeV3
$|\bar{q}qg;1^{-+}\rangle$ $J^{\alpha\beta}_{1^{-+}}$ 5.5 1.811.811.811.812.002.002.002.00 $6.2$ 4040404048484848 $2.00^{+0.13}_{-0.16}$ $0.055^{+0.007}_{-0.008}$ GeV3
$|\bar{q}qg;1^{++}\rangle$ $\tilde{J}^{\alpha\beta}_{1^{++}}$ 5.8 1.841.841.841.842.062.062.062.06 $6.5$ 4040404048484848 $2.11^{+0.17}_{-0.21}$ $0.056^{+0.012}_{-0.013}$ GeV3
$|\bar{q}qg;2^{++}\rangle$ $J^{\alpha_{1}\beta_{1},\alpha_{2}\beta_{2}}_{2^{++}}$ 8.6 3.113.113.113.113.373.373.373.37 $9.5$ 4040404046464646 $2.44^{+0.20}_{-0.24}$
$|\bar{q}qg;2^{-+}\rangle$ $\tilde{J}^{\alpha_{1}\beta_{1},\alpha_{2}\beta_{2}}_{2^{-+}}$ 12.7 2.542.542.542.543.603.603.603.60 $14.0$ 4040404054545454 $3.68^{+0.62}_{-0.18}$
$|\bar{q}qg;2^{-+}\rangle$ $J^{\alpha_{1}\beta_{1},\alpha_{2}\beta_{2}}_{2^{-+}}$ 8.3 3.073.073.073.073.413.413.413.41 $9.5$ 4040404048484848 $2.40^{+0.21}_{-0.25}$
$|\bar{q}qg;2^{++}\rangle$ $\tilde{J}^{\alpha_{1}\beta_{1},\alpha_{2}\beta_{2}}_{2^{++}}$ 11.7 2.472.472.472.473.703.703.703.70 $14.0$ 4040404063636363 $3.46^{+0.27}_{-0.11}$
Table 2: QCD sum rule results for the single-gluon hybrid states $|\bar{s}sg;J^{PC}\rangle$, extracted from the single-gluon hybrid currents given in Eqs. (7-14) and Eqs. (19-30) with the quark-gluon contents $\bar{s}sg$.
 State [$J^{PC}$]     Current   $s_{0}^{min}~{}[{\rm GeV}^{2}]$ Working Regions   Pole [%]   Mass [GeV]    Decay Constant
  $M_{B}^{2}~{}[{\rm GeV}^{2}]$   $s_{0}~{}[{\rm GeV}^{2}]$
$|\bar{s}sg;1^{--}\rangle$ $J^{\alpha\beta}_{1^{--}}$ 4.3 2.072.072.072.072.802.802.802.80 $6.5$ 4040404063636363 $1.94^{+0.20}_{-0.21}$ $0.054^{+0.013}_{-0.016}$ GeV3
$|\bar{s}sg;1^{+-}\rangle$ $\tilde{J}^{\alpha\beta}_{1^{+-}}$ 16.2 3.603.603.603.605.405.405.405.40 $20.0$ 4040404065656565 $4.06^{+0.26}_{-0.16}$ $0.071^{+0.019}_{-0.020}$ GeV3
$|\bar{s}sg;1^{+-}\rangle$ $J^{\alpha\beta}_{1^{+-}}$ 5.9 2.542.542.542.542.722.722.722.72 $6.5$ 4040404045454545 $2.01^{+0.17}_{-0.20}$ $0.050^{+0.005}_{-0.006}$ GeV3
$|\bar{s}sg;1^{--}\rangle$ $\tilde{J}^{\alpha\beta}_{1^{--}}$ 16.9 3.733.733.733.735.305.305.305.30 $20.0$ 4040404061616161 $4.12^{+0.26}_{-0.13}$ $0.070^{+0.019}_{-0.020}$ GeV3
$|\bar{s}sg;0^{++}\rangle$ $J^{\mu\rightarrow 0}_{1^{-+}}$ 20.7 5.185.185.185.187.357.357.357.35 $26.0$ 4040404063636363 $4.50^{+0.23}_{-0.22}$ $0.136^{+0.030}_{-0.034}$ GeV3
$|\bar{s}sg;0^{-+}\rangle$ $\tilde{J}^{\mu\rightarrow 0}_{1^{++}}$ 7.2 3.453.453.453.454.084.084.084.08 $9.5$ 4040404053535353 $2.26^{+0.21}_{-0.24}$ $0.107^{+0.007}_{-0.005}$ GeV3
$|\bar{s}sg;0^{--}\rangle$ $J^{\mu\rightarrow 0}_{1^{+-}}$ 21.6 5.365.365.365.367.237.237.237.23 $26.0$ 4040404059595959 $4.57^{+0.22}_{-0.19}$ $0.134^{+0.031}_{-0.035}$ GeV3
$|\bar{s}sg;0^{+-}\rangle$ $\tilde{J}^{\mu\rightarrow 0}_{1^{--}}$ 7.5 3.413.413.413.413.983.983.983.98 $9.5$ 4040404052525252 $2.30^{+0.20}_{-0.24}$ $0.101^{+0.007}_{-0.006}$ GeV3
$|\bar{s}sg;1^{-+}\rangle$ $J^{\mu}_{1^{-+}}$ 5.1 2.262.262.262.262.542.542.542.54 $6.2$ 4040404049494949 $1.84^{+0.14}_{-0.15}$ $0.300^{+0.063}_{-0.058}$ GeV4
$|\bar{s}sg;1^{++}\rangle$ $\tilde{J}^{\mu}_{1^{++}}$ 14.1 3.643.643.643.644.804.804.804.80 $17.0$ 4040404058585858 $3.65^{+0.17}_{-0.17}$ $1.678^{+0.530}_{-0.502}$ GeV4
$|\bar{s}sg;1^{+-}\rangle$ $J^{\mu}_{1^{+-}}$ 3.9 1.851.851.851.852.432.432.432.43 $6.0$ 4040404062626262 $1.82^{+0.13}_{-0.15}$ $0.278^{+0.059}_{-0.056}$ GeV4
$|\bar{s}sg;1^{--}\rangle$ $\tilde{J}^{\mu}_{1^{--}}$ 13.8 3.503.503.503.504.804.804.804.80 $17.0$ 4040404061616161 $3.64^{+0.17}_{-0.17}$ $1.662^{+0.526}_{-0.498}$ GeV4
$|\bar{s}sg;0^{++}\rangle$ $J_{0^{++}}$ 11.5 3.533.533.533.534.334.334.334.33 $14.0$ 4040404055555555 $3.11^{+0.22}_{-0.27}$ $3.535^{+1.338}_{-1.242}$ GeV4
$|\bar{s}sg;0^{-+}\rangle$ $J_{0^{-+}}$ 11.3 3.513.513.513.514.364.364.364.36 $14.0$ 4040404056565656 $3.08^{+0.23}_{-0.28}$ $3.509^{+1.328}_{-1.233}$ GeV4
$|\bar{s}sg;1^{++}\rangle$ $J^{\alpha\beta}_{1^{++}}$ 6.6 1.951.951.951.952.272.272.272.27 $7.5$ 4040404051515151 $2.34^{+0.14}_{-0.16}$ $0.061^{+0.012}_{-0.014}$ GeV3
$|\bar{s}sg;1^{-+}\rangle$ $\tilde{J}^{\alpha\beta}_{1^{-+}}$ 5.5 1.821.821.821.822.252.252.252.25 $7.0$ 4040404057575757 $2.08^{+0.18}_{-0.24}$ $0.061^{+0.010}_{-0.010}$ GeV3
$|\bar{s}sg;1^{-+}\rangle$ $J^{\alpha\beta}_{1^{-+}}$ 5.5 1.821.821.821.822.252.252.252.25 $7.0$ 4040404057575757 $2.08^{+0.18}_{-0.24}$ $0.061^{+0.010}_{-0.010}$ GeV3
$|\bar{s}sg;1^{++}\rangle$ $\tilde{J}^{\alpha\beta}_{1^{++}}$ 6.6 1.951.951.951.952.272.272.272.27 $7.5$ 4040404051515151 $2.34^{+0.14}_{-0.16}$ $0.061^{+0.012}_{-0.014}$ GeV3
$|\bar{s}sg;2^{++}\rangle$ $J^{\alpha_{1}\beta_{1},\alpha_{2}\beta_{2}}_{2^{++}}$ 9.2 3.223.223.223.223.603.603.603.60 $10.5$ 4040404049494949 $2.59^{+0.19}_{-0.23}$
$|\bar{s}sg;2^{-+}\rangle$ $\tilde{J}^{\alpha_{1}\beta_{1},\alpha_{2}\beta_{2}}_{2^{-+}}$ 13.4 2.552.552.552.554.294.294.294.29 $16.0$ 4040404066666666 $3.72^{+0.72}_{-0.13}$
$|\bar{s}sg;2^{-+}\rangle$ $J^{\alpha_{1}\beta_{1},\alpha_{2}\beta_{2}}_{2^{-+}}$ 8.1 3.043.043.043.043.723.723.723.72 $10.5$ 4040404056565656 $2.51^{+0.20}_{-0.24}$
$|\bar{s}sg;2^{++}\rangle$ $\tilde{J}^{\alpha_{1}\beta_{1},\alpha_{2}\beta_{2}}_{2^{++}}$ 11.8 2.362.362.362.364.474.474.474.47 $16.0$ 4040404078787878 $3.54^{+0.42}_{-0.16}$
(a) (a)
(b) (b)
(c) (c)
Figure 5: Decay mechanisms of the single-gluon hybrid states through (a) the normal process with one quark-antiquark pair excited from the valence gluon, and (b,c) the abnormal processes with the $\eta^{(\prime)}$ produced by the QCD axial anomaly.
Table 3: Partial decay widths of the hybrid states $|\bar{q}qg;1^{-}1^{-+}\rangle$, $|\bar{q}qg;0^{+}1^{-+}\rangle$, and $|\bar{s}sg;0^{+}1^{-+}\rangle$, in units of MeV. $\Gamma(\pi_{1}/\eta_{1}\xrightarrow{a,b,c}M_{1}M_{2})$ are related to the processes depicted in Figs. 5(a,b,c), respectively. We simply sum over the partial decay widths to obtain the total decay widths, as listed in the last row. Note that the decay channels $\pi_{1}\xrightarrow{b,c}\eta\pi/\eta^{\prime}\pi$ and $\pi_{1}/\eta_{1}\rightarrow K^{*}(892)\bar{K}/K_{1}(1270)\bar{K}/K^{*}(892)\bar{K}^{*}(892)$ have not been investigated in our previous QCD sum rule studies Chen:2010ic ; Huang:2010dc ; Chen:2022qpd .
              Channel        $|\bar{q}qg;1^{-}1^{-+}\rangle$        $|\bar{q}qg;0^{+}1^{-+}\rangle$        $|\bar{s}sg;0^{+}1^{-+}\rangle$
$M=1.67^{+0.15}_{-0.17}$ GeV $M=1.67^{+0.15}_{-0.17}$ GeV $M=1.84^{+0.14}_{-0.15}$ GeV
$\pi_{1}/\eta_{1}\rightarrow\rho\pi$ $242^{+310}_{-179}$
$\pi_{1}/\eta_{1}\rightarrow b_{1}(1235)\pi$ $14.5^{+25.9}_{-13.9}$
$\pi_{1}/\eta_{1}\rightarrow f_{1}(1285)\pi$ $35.9^{+53.9}_{-36.4}$
$\pi_{1}/\eta_{1}\rightarrow\eta\pi$ $2.3^{+2.5}_{-1.2}$
[…]
$\pi_{1}/\eta_{1}\rightarrow\eta^{\prime}\pi$ $0.43^{+0.50}_{-0.28}$
[…]
$\pi_{1}/\eta_{1}\rightarrow a_{1}(1260)\pi$ $79.5^{+112.4}_{-~{}74.9}$
[…]
$\pi_{1}/\eta_{1}\rightarrow K^{*}(892)\bar{K}+c.c.$ $25.3^{+34.7}_{-24.7}$ $25.3^{+34.7}_{-24.7}$ $73.9^{+85.7}_{-58.0}$
$\pi_{1}/\eta_{1}\rightarrow K_{1}(1270)\bar{K}+c.c.$ $14.6^{+19.8}_{-14.6}$
$\pi_{1}/\eta_{1}\rightarrow K^{*}(892)\bar{K}^{*}(892)$ $0.08^{+0.39}_{-0.08}$
Sum $530^{+540}_{-330}$ $120^{+160}_{-110}$ $100^{+110}_{-~{}80}$
In this paper we study the single-gluon hybrid states with various (exotic) quantum numbers. We systematically construct twenty-four single-gluon hybrid currents, and use eighteen of them to perform QCD sum rule analyses. We calculate the masses of forty-four single-gluon hybrid states with the quark-gluon contents $\bar{q}qg$ ($q=u/d$) and $\bar{s}sg$. The obtained results are summarized in Table 1 and Table 2. Especially, the masses and decay constants of the $J^{PC}=1^{-+}$ hybrid states are extracted from the current $J^{\mu}_{1^{-+}}$ given in Eq. (11) to be:
$M_{|\bar{q}qg;1^{-+}\rangle}$ $=$ $1.67^{+0.15}_{-0.17}{\rm~{}GeV}\,,$
$f_{|\bar{q}qg;1^{-+}\rangle}$ $=$ $0.243^{+0.057}_{-0.052}{\rm~{}GeV}\,,$
$M_{|\bar{s}sg;1^{-+}\rangle}$ $=$ $1.84^{+0.14}_{-0.15}{\rm~{}GeV}\,,$
$f_{|\bar{s}sg;1^{-+}\rangle}$ $=$ $0.300^{+0.063}_{-0.058}{\rm~{}GeV}\,.$
Since we do not differentiate the up and down quarks within the QCD sum rule method, the masses and decay constants of the states in the same isospin multiplet are calculated to be the same:
$M_{|\bar{q}qg;1^{-}1^{-+}\rangle}=M_{|\bar{q}qg;0^{+}1^{-+}\rangle}$ $=$ $1.67^{+0.15}_{-0.17}{\rm~{}GeV}\,,$
$f_{|\bar{q}qg;1^{-}1^{-+}\rangle}=f_{|\bar{q}qg;0^{+}1^{-+}\rangle}$ $=$ $0.243^{+0.057}_{-0.052}{\rm~{}GeV}\,,$
$M_{|\bar{s}sg;0^{+}1^{-+}\rangle}$ $=$ $1.84^{+0.14}_{-0.15}{\rm~{}GeV}\,,$
$f_{|\bar{s}sg;0^{+}1^{-+}\rangle}$ $=$ $0.300^{+0.063}_{-0.058}{\rm~{}GeV}\,.$
As shown in Fig. 5(a), a single-gluon hybrid state can decay after exciting one $\bar{q}q/\bar{s}s$ pair from the valence gluon, followed by reorganizing two color-octet $\bar{q}q/\bar{s}s$ pairs into two color-singlet mesons. This decay process has been systematically studied in Refs. Chen:2010ic ; Huang:2010dc for the $J^{PC}=1^{-+}$ hybrid states through the QCD sum rule method, and in the present study we update these calculations. Besides the “normal” decay process depicted in Fig. 5(a), there also exist the “abnormal” decay processes depicted in Figs. 5(b,c), where the $\eta^{(\prime)}$ mesons are produced by the QCD axial anomaly. These abnormal decay processes have been systematically studied in Ref. Chen:2022qpd , and in the present study we also update these calculations. The obtained results are summarized in Table 3, and especially,
$\Gamma_{|\bar{q}qg;1^{-}1^{-+}\rangle}$ $=$ $530^{+540}_{-330}{\rm~{}MeV}\,,$
$\Gamma_{|\bar{q}qg;0^{+}1^{-+}\rangle}$ $=$ $120^{+160}_{-110}{\rm~{}MeV}\,,$
$\Gamma_{|\bar{s}sg;0^{+}1^{-+}\rangle}$ $=$ $100^{+110}_{-~{}80}{\rm~{}MeV}\,.$
The above QCD sum rule results suggest that the $\pi_{1}(1600)$ and $\eta_{1}(1855)$ can be respectively interpreted as the single-gluon hybrid states $|\bar{q}qg;1^{-}1^{-+}\rangle$ and $|\bar{s}sg;0^{+}1^{-+}\rangle$, so there exists another isoscalar state $|\bar{q}qg;0^{+}1^{-+}\rangle$, whose mass and width are smaller than those of $\eta_{1}(1855)$. Considering the uncertainties, our results suggest that the $\pi_{1}(1600)$ and $\eta_{1}(1855)$ may also be respectively interpreted as the single-gluon hybrid states $|\bar{q}qg;1^{-}1^{-+}\rangle$ and $|\bar{q}qg;0^{+}1^{-+}\rangle$, so there exists another isoscalar state $|\bar{s}sg;0^{+}1^{-+}\rangle$, whose mass and width are larger than those of $\eta_{1}(1855)$. To differentiate these two assignments, it is useful to examine the $a_{1}(1260)\pi$ decay channel. We find in Table 3 that the $\eta^{(\prime)}$-relevant decay modes, as the characteristic decay modes of hybrid states, are enhanced to some extent. To verify whether the exotic $\pi_{1}$ and $\eta_{1}$ resonances are hybrid states or not, we propose to examine these decay modes in future BESIII, Belle-II, GlueX, LHCb, and PANDA experiments.
Acknowledgments
This project is supported by the National Natural Science Foundation of China under Grant No. 12075019, the Jiangsu Provincial Double-Innovation Program under Grant No. JSSCRC2021488, and the Fundamental Research Funds for the Central Universities.
Appendix A Spectral densities
In this appendix we show the OPE spectral densities extracted from the single-gluon hybrid currents given in Eqs. (7-14) and Eqs. (19-30) with the quark-gluon content $\bar{s}sg$. Those with the quark-gluon content $\bar{q}qg$ ($q=u/d$) can be easily derived by replacing $m_{s}\rightarrow 0$, $\left\langle{\bar{s}s}\right\rangle\rightarrow\left\langle{\bar{q}q}\right\rangle$, and $\left\langle{{g_{s}}\bar{s}\sigma Gs}\right\rangle\rightarrow\left\langle{{g_{s}}\bar{q}\sigma Gq}\right\rangle$.
$\Pi^{\alpha\beta}_{1^{--}}{\left({M_{B}^{2},{s_{0}}}\right)}$ $=\int_{4m_{s}^{2}}^{{s_{0}}}{(\frac{{{s^{3}}{\alpha_{s}}}}{{240{\pi^{3}}}}}-\frac{{m_{s}^{2}{s^{2}}{\alpha_{s}}}}{{24{\pi^{3}}}}+s(\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle}}{{48{\pi^{2}}}}-\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle{\alpha_{s}}}}{{1152{\pi^{3}}}}-\frac{{2{m_{s}}\left\langle{\bar{s}s}\right\rangle{\alpha_{s}}}}{{9\pi}})$ (57)
$-\frac{{\left\langle{g_{s}^{3}{G^{3}}}\right\rangle}}{{96{\pi^{2}}}}-\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle m_{s}^{2}}}{{24{\pi^{2}}}})\times{e^{-s/M_{B}^{2}}}ds+(-\frac{{{{\left\langle{{\alpha_{s}}GG}\right\rangle}^{2}}}}{{4608{\pi^{2}}}}-\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle{m_{s}}\left\langle{\bar{s}s}\right\rangle}}{{18}}$
$-\frac{4}{9}m_{s}^{2}\pi{\left\langle{\bar{s}s}\right\rangle^{2}}{\alpha_{s}}+\frac{8}{9}\pi\left\langle{\bar{s}s}\right\rangle\left\langle{{g_{s}}\bar{s}\sigma Gs}\right\rangle{\alpha_{s}})\,,$
$\widetilde{\Pi}^{\alpha\beta}_{1^{+-}}{\left({M_{B}^{2},{s_{0}}}\right)}$ $=\int_{4m_{s}^{2}}^{{s_{0}}}{(\frac{{{s^{3}}{\alpha_{s}}}}{{60{\pi^{3}}}}}-\frac{{m_{s}^{2}{s^{2}}{\alpha_{s}}}}{{6{\pi^{3}}}}+s(-\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle}}{{12{\pi^{2}}}}-\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle{\alpha_{s}}}}{{288{\pi^{3}}}}-\frac{{8{m_{s}}\left\langle{\bar{s}s}\right\rangle{\alpha_{s}}}}{{9\pi}})$ (58)
$+\frac{{\left\langle{g_{s}^{3}{G^{3}}}\right\rangle}}{{24{\pi^{2}}}}+\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle m_{s}^{2}}}{{6{\pi^{2}}}})\times{e^{-s/M_{B}^{2}}}ds+(\frac{{{{\left\langle{{\alpha_{s}}GG}\right\rangle}^{2}}}}{{1152{\pi^{2}}}}-\frac{{\left\langle{g_{s}^{3}{G^{3}}}\right\rangle m_{s}^{2}}}{{12{\pi^{2}}}}$
$+\frac{2}{9}\left\langle{{\alpha_{s}}GG}\right\rangle{m_{s}}\left\langle{\bar{s}s}\right\rangle-\frac{{16}}{9}\pi m_{s}^{2}{\left\langle{\bar{s}s}\right\rangle^{2}}{\alpha_{s}}+\frac{{32}}{9}\pi\left\langle{\bar{s}s}\right\rangle\left\langle{{g_{s}}\bar{s}\sigma Gs}\right\rangle{\alpha_{s}})\,,$
$\Pi^{\alpha\beta}_{1^{+-}}{\left({M_{B}^{2},{s_{0}}}\right)}$ $=\int_{4m_{s}^{2}}^{{s_{0}}}{(\frac{{{s^{3}}{\alpha_{s}}}}{{240{\pi^{3}}}}}-\frac{{m_{s}^{2}{s^{2}}{\alpha_{s}}}}{{8{\pi^{3}}}}+s(\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle}}{{48{\pi^{2}}}}-\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle{\alpha_{s}}}}{{1152{\pi^{3}}}}+\frac{{2{m_{s}}\left\langle{\bar{s}s}\right\rangle{\alpha_{s}}}}{{3\pi}})$ (59)
$-\frac{{\left\langle{g_{s}^{3}{G^{3}}}\right\rangle}}{{96{\pi^{2}}}}-\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle m_{s}^{2}}}{{8{\pi^{2}}}})\times{e^{-s/M_{B}^{2}}}ds+(-\frac{{{{\left\langle{{\alpha_{s}}GG}\right\rangle}^{2}}}}{{4608{\pi^{2}}}}+\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle{m_{s}}\left\langle{\bar{s}s}\right\rangle}}{6}$
$-\frac{4}{9}m_{s}^{2}\pi{\left\langle{\bar{s}s}\right\rangle^{2}}{\alpha_{s}}-\frac{8}{9}\pi\left\langle{\bar{s}s}\right\rangle\left\langle{{g_{s}}\bar{s}\sigma Gs}\right\rangle{\alpha_{s}})\,,$
$\widetilde{\Pi}^{\alpha\beta}_{1^{--}}{\left({M_{B}^{2},{s_{0}}}\right)}$ $=\int_{4m_{s}^{2}}^{{s_{0}}}{(\frac{{{s^{3}}{\alpha_{s}}}}{{60{\pi^{3}}}}}-\frac{{m_{s}^{2}{s^{2}}{\alpha_{s}}}}{{2{\pi^{3}}}}+s(-\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle}}{{12{\pi^{2}}}}-\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle{\alpha_{s}}}}{{288{\pi^{3}}}}+\frac{{8{m_{s}}\left\langle{\bar{s}s}\right\rangle{\alpha_{s}}}}{{3\pi}})$ (60)
$+\frac{{\left\langle{g_{s}^{3}{G^{3}}}\right\rangle}}{{24{\pi^{2}}}}+\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle m_{s}^{2}}}{{2{\pi^{2}}}})\times{e^{-s/M_{B}^{2}}}ds+(\frac{{{{\left\langle{{\alpha_{s}}GG}\right\rangle}^{2}}}}{{1152{\pi^{2}}}}-\frac{{\left\langle{g_{s}^{3}{G^{3}}}\right\rangle m_{s}^{2}}}{{4{\pi^{2}}}}$
$-\frac{2}{3}\left\langle{{\alpha_{s}}GG}\right\rangle{m_{s}}\left\langle{\bar{s}s}\right\rangle-\frac{{16}}{9}\pi m_{s}^{2}{\left\langle{\bar{s}s}\right\rangle^{2}}{\alpha_{s}}-\frac{{32}}{9}\pi\left\langle{\bar{s}s}\right\rangle\left\langle{{g_{s}}\bar{s}\sigma Gs}\right\rangle{\alpha_{s}})\,,$
$\Pi^{\mu}_{1^{-+}}{\left({M_{B}^{2},{s_{0}}}\right)}$ $=\int_{4m_{s}^{2}}^{{s_{0}}}{(\frac{{{s^{3}}{\alpha_{s}}}}{{60{\pi^{3}}}}}-\frac{{m_{s}^{2}{s^{2}}{\alpha_{s}}}}{{3{\pi^{3}}}}+s(\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle}}{{36{\pi^{2}}}}+\frac{{13\left\langle{{\alpha_{s}}GG}\right\rangle{\alpha_{s}}}}{{432{\pi^{3}}}}+\frac{{8{m_{s}}\left\langle{\bar{s}s}\right\rangle{\alpha_{s}}}}{{9\pi}})$ (61)
$+\frac{{\left\langle{g_{s}^{3}{G^{3}}}\right\rangle}}{{32{\pi^{2}}}}-\frac{{3\left\langle{{\alpha_{s}}GG}\right\rangle m_{s}^{2}{\alpha_{s}}}}{{64{\pi^{3}}}}-\frac{{3{m_{s}}\left\langle{{g_{s}}\bar{s}\sigma Gs}\right\rangle{\alpha_{s}}}}{{4\pi}})\times{e^{-s/M_{B}^{2}}}ds$
$+(\frac{{{{\left\langle{{\alpha_{s}}GG}\right\rangle}^{2}}}}{{3456{\pi^{2}}}}-\frac{{\left\langle{g_{s}^{3}{G^{3}}}\right\rangle m_{s}^{2}}}{{16{\pi^{2}}}}-\frac{2}{9}\left\langle{{\alpha_{s}}GG}\right\rangle{m_{s}}\left\langle{\bar{s}s}\right\rangle+\frac{{11}}{9}\pi\left\langle{\bar{s}s}\right\rangle\left\langle{{g_{s}}\bar{s}\sigma Gs}\right\rangle{\alpha_{s}})\,,$
$\widetilde{\Pi}^{\mu}_{1^{++}}{\left({M_{B}^{2},{s_{0}}}\right)}$ $=\int_{4m_{s}^{2}}^{{s_{0}}}{(\frac{{{s^{3}}{\alpha_{s}}}}{{15{\pi^{3}}}}}-\frac{{4m_{s}^{2}{s^{2}}{\alpha_{s}}}}{{3{\pi^{3}}}}+s(-\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle}}{{9{\pi^{2}}}}-\frac{{5\left\langle{{\alpha_{s}}GG}\right\rangle{\alpha_{s}}}}{{108{\pi^{3}}}}+\frac{{32{m_{s}}\left\langle{\bar{s}s}\right\rangle{\alpha_{s}}}}{{9\pi}})$ (62)
$-\frac{{\left\langle{g_{s}^{3}{G^{3}}}\right\rangle}}{{4{\pi^{2}}}}+\frac{{3\left\langle{{\alpha_{s}}GG}\right\rangle m_{s}^{2}{\alpha_{s}}}}{{16{\pi^{3}}}}+\frac{{3{m_{s}}\left\langle{{g_{s}}\bar{s}\sigma Gs}\right\rangle{\alpha_{s}}}}{\pi})\times{e^{-s/M_{B}^{2}}}ds+(-\frac{{{{\left\langle{{\alpha_{s}}GG}\right\rangle}^{2}}}}{{864{\pi^{2}}}}$
$+\frac{{\left\langle{g_{s}^{3}{G^{3}}}\right\rangle m_{s}^{2}}}{{3{\pi^{2}}}}+\frac{8}{9}\left\langle{{\alpha_{s}}GG}\right\rangle{m_{s}}\left\langle{\bar{s}s}\right\rangle-\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle{m_{s}}\left\langle{\bar{s}s}\right\rangle{\alpha_{s}}}}{{3\pi}}-\frac{{28}}{9}\pi\left\langle{\bar{s}s}\right\rangle\left\langle{{g_{s}}\bar{s}\sigma Gs}\right\rangle{\alpha_{s}})\,,$
$\Pi^{\mu}_{1^{+-}}{\left({M_{B}^{2},{s_{0}}}\right)}$ $=\int_{4m_{s}^{2}}^{{s_{0}}}{(\frac{{{s^{3}}{\alpha_{s}}}}{{60{\pi^{3}}}}}-\frac{{5m_{s}^{2}{s^{2}}{\alpha_{s}}}}{{12{\pi^{3}}}}+s(\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle}}{{36{\pi^{2}}}}+\frac{{13\left\langle{{\alpha_{s}}GG}\right\rangle{\alpha_{s}}}}{{432{\pi^{3}}}}+\frac{{16{m_{s}}\left\langle{\bar{s}s}\right\rangle{\alpha_{s}}}}{{9\pi}})$ (63)
$+\frac{{\left\langle{g_{s}^{3}{G^{3}}}\right\rangle}}{{32{\pi^{2}}}}-\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle m_{s}^{2}}}{{4{\pi^{2}}}}-\frac{{3\left\langle{{\alpha_{s}}GG}\right\rangle m_{s}^{2}{\alpha_{s}}}}{{64{\pi^{3}}}}+\frac{{3{m_{s}}\left\langle{{g_{s}}\bar{s}\sigma Gs}\right\rangle{\alpha_{s}}}}{{4\pi}})\times{e^{-s/M_{B}^{2}}}ds$
$+(\frac{{{{\left\langle{{\alpha_{s}}GG}\right\rangle}^{2}}}}{{3456{\pi^{2}}}}-\frac{{\left\langle{g_{s}^{3}{G^{3}}}\right\rangle m_{s}^{2}}}{{48{\pi^{2}}}}+\frac{4}{9}\left\langle{{\alpha_{s}}GG}\right\rangle{m_{s}}\left\langle{\bar{s}s}\right\rangle-\frac{{11}}{9}\pi\left\langle{\bar{s}s}\right\rangle\left\langle{{g_{s}}\bar{s}\sigma Gs}\right\rangle{\alpha_{s}})\,,$
$\widetilde{\Pi}^{\mu}_{1^{--}}{\left({M_{B}^{2},{s_{0}}}\right)}$ $=\int_{4m_{s}^{2}}^{{s_{0}}}{(\frac{{{s^{3}}{\alpha_{s}}}}{{15{\pi^{3}}}}}-\frac{{5m_{s}^{2}{s^{2}}{\alpha_{s}}}}{{3{\pi^{3}}}}+s(-\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle}}{{9{\pi^{2}}}}-\frac{{5\left\langle{{\alpha_{s}}GG}\right\rangle{\alpha_{s}}}}{{108{\pi^{3}}}}+\frac{{64{m_{s}}\left\langle{\bar{s}s}\right\rangle{\alpha_{s}}}}{{9\pi}})$ (64)
$-\frac{{\left\langle{g_{s}^{3}{G^{3}}}\right\rangle}}{{4{\pi^{2}}}}+\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle m_{s}^{2}}}{{{\pi^{2}}}}+\frac{{3\left\langle{{\alpha_{s}}GG}\right\rangle m_{s}^{2}{\alpha_{s}}}}{{16{\pi^{3}}}}-\frac{{3{m_{s}}\left\langle{{g_{s}}\bar{s}\sigma Gs}\right\rangle{\alpha_{s}}}}{\pi})\times{e^{-s/M_{B}^{2}}}ds$
$+(-\frac{{{{\left\langle{{\alpha_{s}}GG}\right\rangle}^{2}}}}{{864{\pi^{2}}}}-\frac{{16}}{9}\left\langle{{\alpha_{s}}GG}\right\rangle{m_{s}}\left\langle{\bar{s}s}\right\rangle-\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle{m_{s}}\left\langle{\bar{s}s}\right\rangle{\alpha_{s}}}}{{3\pi}}+\frac{{28}}{9}\pi\left\langle{\bar{s}s}\right\rangle\left\langle{{g_{s}}\bar{s}\sigma Gs}\right\rangle{\alpha_{s}})\,,$
$\Pi^{\mu}_{0^{++}}{\left({M_{B}^{2},{s_{0}}}\right)}$ $=\int_{4m_{s}^{2}}^{{s_{0}}}{(\frac{{{s^{3}}{\alpha_{s}}}}{{120{\pi^{3}}}}}+s(-\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle}}{{24{\pi^{2}}}}+\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle{\alpha_{s}}}}{{576{\pi^{3}}}}-\frac{{4{m_{s}}\left\langle{\bar{s}s}\right\rangle{\alpha_{s}}}}{{3\pi}})-\frac{{3\left\langle{g_{s}^{3}{G^{3}}}\right\rangle}}{{32{\pi^{2}}}}$ (65)
$+\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle m_{s}^{2}}}{{12{\pi^{2}}}}-\frac{{3\left\langle{{\alpha_{s}}GG}\right\rangle m_{s}^{2}{\alpha_{s}}}}{{64{\pi^{3}}}}+\frac{{32}}{9}\pi{\left\langle{\bar{s}s}\right\rangle^{2}}{\alpha_{s}}+\frac{{13\left\langle{{g_{s}}\bar{s}\sigma Gs}\right\rangle{m_{s}}{\alpha_{s}}}}{{12\pi}})\times{e^{-s/M_{B}^{2}}}ds$
$+(-\frac{{{{\left\langle{{\alpha_{s}}GG}\right\rangle}^{2}}}}{{2304{\pi^{2}}}}-\frac{{\left\langle{g_{s}^{3}{G^{3}}}\right\rangle m_{s}^{2}}}{{16{\pi^{2}}}}+\frac{1}{3}\left\langle{{\alpha_{s}}GG}\right\rangle{m_{s}}\left\langle{\bar{s}s}\right\rangle-\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle{m_{s}}\left\langle{\bar{s}s}\right\rangle{\alpha_{s}}}}{{8\pi}}$
$-\frac{{8\pi m_{s}^{2}{{\left\langle{\bar{s}s}\right\rangle}^{2}}{\alpha_{s}}}}{3}-\frac{1}{3}\pi\left\langle{\bar{s}s}\right\rangle\left\langle{{g_{s}}\bar{s}\sigma Gs}\right\rangle{\alpha_{s}})\,,$
$\widetilde{\Pi}^{\mu}_{0^{-+}}{\left({M_{B}^{2},{s_{0}}}\right)}$ $=\int_{4m_{s}^{2}}^{{s_{0}}}{(\frac{{{s^{3}}{\alpha_{s}}}}{{30{\pi^{3}}}}}+s(\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle}}{{6{\pi^{2}}}}+\frac{{37\left\langle{{\alpha_{s}}GG}\right\rangle{\alpha_{s}}}}{{144{\pi^{3}}}}-\frac{{16{m_{s}}\left\langle{\bar{s}s}\right\rangle{\alpha_{s}}}}{{3\pi}})+\frac{{\left\langle{g_{s}^{3}{G^{3}}}\right\rangle}}{{4{\pi^{2}}}}-\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle m_{s}^{2}}}{{3{\pi^{2}}}}$ (66)
$-\frac{{13\left\langle{{\alpha_{s}}GG}\right\rangle m_{s}^{2}{\alpha_{s}}}}{{16{\pi^{3}}}}+\frac{{128}}{9}\pi{\left\langle{\bar{s}s}\right\rangle^{2}}{\alpha_{s}}-\frac{{5\left\langle{{g_{s}}\bar{s}\sigma Gs}\right\rangle{m_{s}}{\alpha_{s}}}}{{3\pi}})\times{e^{-s/M_{B}^{2}}}ds$
$+(\frac{{{{\left\langle{{\alpha_{s}}GG}\right\rangle}^{2}}}}{{576{\pi^{2}}}}-\frac{{\left\langle{g_{s}^{3}{G^{3}}}\right\rangle m_{s}^{2}}}{{2{\pi^{2}}}}-\frac{4}{3}\left\langle{{\alpha_{s}}GG}\right\rangle{m_{s}}\left\langle{\bar{s}s}\right\rangle-\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle{m_{s}}\left\langle{\bar{s}s}\right\rangle{\alpha_{s}}}}{{2\pi}}$
$-\frac{{32\pi m_{s}^{2}{{\left\langle{\bar{s}s}\right\rangle}^{2}}{\alpha_{s}}}}{3}+\frac{{20}}{3}\pi\left\langle{\bar{s}s}\right\rangle\left\langle{{g_{s}}\bar{s}\sigma Gs}\right\rangle{\alpha_{s}})\,,$
$\Pi^{\mu}_{0^{--}}{\left({M_{B}^{2},{s_{0}}}\right)}$ $=\int_{4m_{s}^{2}}^{{s_{0}}}{(\frac{{{s^{3}}{\alpha_{s}}}}{{120{\pi^{3}}}}}-\frac{{m_{s}^{2}s^{2}{\alpha_{s}}}}{{4{\pi^{3}}}}+s(-\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle}}{{24{\pi^{2}}}}+\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle{\alpha_{s}}}}{{576{\pi^{3}}}}+\frac{{4{m_{s}}\left\langle{\bar{s}s}\right\rangle{\alpha_{s}}}}{{3\pi}})-\frac{{3\left\langle{g_{s}^{3}{G^{3}}}\right\rangle}}{{32{\pi^{2}}}}$ (67)
$+\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle m_{s}^{2}}}{{3{\pi^{2}}}}-\frac{{3\left\langle{{\alpha_{s}}GG}\right\rangle m_{s}^{2}{\alpha_{s}}}}{{64{\pi^{3}}}}-\frac{{32}}{9}\pi{\left\langle{\bar{s}s}\right\rangle^{2}}{\alpha_{s}}-\frac{{13\left\langle{{g_{s}}\bar{s}\sigma Gs}\right\rangle{m_{s}}{\alpha_{s}}}}{{12\pi}})\times{e^{-s/M_{B}^{2}}}ds$
$+(-\frac{{{{\left\langle{{\alpha_{s}}GG}\right\rangle}^{2}}}}{{2304{\pi^{2}}}}-\frac{{3\left\langle{g_{s}^{3}{G^{3}}}\right\rangle m_{s}^{2}}}{{16{\pi^{2}}}}-\frac{1}{3}\left\langle{{\alpha_{s}}GG}\right\rangle{m_{s}}\left\langle{\bar{s}s}\right\rangle-\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle{m_{s}}\left\langle{\bar{s}s}\right\rangle{\alpha_{s}}}}{{8\pi}}$
$-\frac{{8\pi m_{s}^{2}{{\left\langle{\bar{s}s}\right\rangle}^{2}}{\alpha_{s}}}}{3}+\frac{1}{3}\pi\left\langle{\bar{s}s}\right\rangle\left\langle{{g_{s}}\bar{s}\sigma Gs}\right\rangle{\alpha_{s}})\,,$
$\widetilde{\Pi}^{\mu}_{0^{+-}}{\left({M_{B}^{2},{s_{0}}}\right)}$ $=\int_{4m_{s}^{2}}^{{s_{0}}}{(\frac{{{s^{3}}{\alpha_{s}}}}{{30{\pi^{3}}}}}-\frac{{m_{s}^{2}s^{2}{\alpha_{s}}}}{{{\pi^{3}}}}+s(+\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle}}{{6{\pi^{2}}}}+\frac{{37\left\langle{{\alpha_{s}}GG}\right\rangle{\alpha_{s}}}}{{144{\pi^{3}}}}+\frac{{16{m_{s}}\left\langle{\bar{s}s}\right\rangle{\alpha_{s}}}}{{3\pi}})+\frac{{\left\langle{g_{s}^{3}{G^{3}}}\right\rangle}}{{4{\pi^{2}}}}$ (68)
$-\frac{{4\left\langle{{\alpha_{s}}GG}\right\rangle m_{s}^{2}}}{{3{\pi^{2}}}}-\frac{{13\left\langle{{\alpha_{s}}GG}\right\rangle m_{s}^{2}{\alpha_{s}}}}{{16{\pi^{3}}}}-\frac{{128}}{9}\pi{\left\langle{\bar{s}s}\right\rangle^{2}}{\alpha_{s}}+\frac{{5\left\langle{{g_{s}}\bar{s}\sigma Gs}\right\rangle{m_{s}}{\alpha_{s}}}}{{3\pi}})\times{e^{-s/M_{B}^{2}}}ds$
$+(\frac{{{{\left\langle{{\alpha_{s}}GG}\right\rangle}^{2}}}}{{576{\pi^{2}}}}-\frac{{\left\langle{g_{s}^{3}{G^{3}}}\right\rangle m_{s}^{2}}}{{2{\pi^{2}}}}+\frac{4}{3}\left\langle{{\alpha_{s}}GG}\right\rangle{m_{s}}\left\langle{\bar{s}s}\right\rangle-\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle{m_{s}}\left\langle{\bar{s}s}\right\rangle{\alpha_{s}}}}{{2\pi}}$
$-\frac{{32\pi m_{s}^{2}{{\left\langle{\bar{s}s}\right\rangle}^{2}}{\alpha_{s}}}}{3}-\frac{{20}}{3}\pi\left\langle{\bar{s}s}\right\rangle\left\langle{{g_{s}}\bar{s}\sigma Gs}\right\rangle{\alpha_{s}})\,,$
$\Pi_{0^{++}}{\left({M_{B}^{2},{s_{0}}}\right)}$ $=\int_{4m_{s}^{2}}^{{s_{0}}}{(\frac{{{s^{3}}{\alpha_{s}}}}{{6{\pi^{3}}}}}-\frac{{4m_{s}^{2}{s^{2}}{\alpha_{s}}}}{{{\pi^{3}}}}+s(\frac{{37\left\langle{{\alpha_{s}}GG}\right\rangle{\alpha_{s}}}}{{48{\pi^{3}}}}+\frac{{16{m_{s}}\left\langle{\bar{s}s}\right\rangle{\alpha_{s}}}}{\pi})-\frac{{\left\langle{g_{s}^{3}{G^{3}}}\right\rangle}}{{4{\pi^{2}}}}$ (69)
$-\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle m_{s}^{2}}}{{{\pi^{2}}}}-\frac{{3\left\langle{{\alpha_{s}}GG}\right\rangle m_{s}^{2}{\alpha_{s}}}}{{2{\pi^{3}}}}+\frac{{12{m_{s}}\left\langle{{g_{s}}\bar{s}\sigma Gs}\right\rangle{\alpha_{s}}}}{\pi})\times{e^{-s/M_{B}^{2}}}ds+(\frac{{3\left\langle{g_{s}^{3}{G^{3}}}\right\rangle m_{s}^{2}}}{{4{\pi^{2}}}}$
$+\frac{8}{3}\left\langle{{\alpha_{s}}GG}\right\rangle{m_{s}}\left\langle{\bar{s}s}\right\rangle+\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle{m_{s}}\left\langle{\bar{s}s}\right\rangle{\alpha_{s}}}}{\pi}+\frac{{32\pi m_{s}^{2}{{\left\langle{\bar{s}s}\right\rangle}^{2}}{\alpha_{s}}}}{3}-16\pi\left\langle{\bar{s}s}\right\rangle\left\langle{{g_{s}}\bar{s}\sigma Gs}\right\rangle{\alpha_{s}})\,,$
$\Pi_{0^{-+}}{\left({M_{B}^{2},{s_{0}}}\right)}$ $=\int_{4m_{s}^{2}}^{{s_{0}}}{(\frac{{{s^{3}}{\alpha_{s}}}}{{6{\pi^{3}}}}}-\frac{{4m_{s}^{2}{s^{2}}{\alpha_{s}}}}{{{\pi^{3}}}}+s(\frac{{37\left\langle{{\alpha_{s}}GG}\right\rangle{\alpha_{s}}}}{{48{\pi^{3}}}}+\frac{{16{m_{s}}\left\langle{\bar{s}s}\right\rangle{\alpha_{s}}}}{\pi})-\frac{{\left\langle{g_{s}^{3}{G^{3}}}\right\rangle}}{{4{\pi^{2}}}}+\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle m_{s}^{2}}}{{{\pi^{2}}}}$ (70)
$-\frac{{3\left\langle{{\alpha_{s}}GG}\right\rangle m_{s}^{2}{\alpha_{s}}}}{{2{\pi^{3}}}}-\frac{{12{m_{s}}\left\langle{{g_{s}}\bar{s}\sigma Gs}\right\rangle{\alpha_{s}}}}{\pi})\times{e^{-s/M_{B}^{2}}}ds+(\frac{{\left\langle{g_{s}^{3}{G^{3}}}\right\rangle m_{s}^{2}}}{{4{\pi^{2}}}}-\frac{8}{3}\left\langle{{\alpha_{s}}GG}\right\rangle{m_{s}}\left\langle{\bar{s}s}\right\rangle$
$+\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle{m_{s}}\left\langle{\bar{s}s}\right\rangle{\alpha_{s}}}}{\pi}+\frac{{32\pi m_{s}^{2}{{\left\langle{\bar{s}s}\right\rangle}^{2}}{\alpha_{s}}}}{3}+16\pi\left\langle{\bar{s}s}\right\rangle\left\langle{{g_{s}}\bar{s}\sigma Gs}\right\rangle{\alpha_{s}})\,,$
$\Pi^{\alpha\beta}_{1^{++}}{\left({M_{B}^{2},{s_{0}}}\right)}$ $=\int_{4m_{s}^{2}}^{{s_{0}}}{(\frac{{{s^{3}}{\alpha_{s}}}}{{60{\pi^{3}}}}}-\frac{{m_{s}^{2}{s^{2}}{\alpha_{s}}}}{{3{\pi^{3}}}}+s(\frac{{19\left\langle{{\alpha_{s}}GG}\right\rangle{\alpha_{s}}}}{{864{\pi^{3}}}}+\frac{{8{m_{s}}\left\langle{\bar{s}s}\right\rangle{\alpha_{s}}}}{{9\pi}})-\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle m_{s}^{2}}}{{6{\pi^{2}}}}$ (71)
$+\frac{{{m_{s}}\left\langle{{g_{s}}\bar{s}\sigma Gs}\right\rangle{\alpha_{s}}}}{\pi})\times{e^{-s/M_{B}^{2}}}ds+(-\frac{{\left\langle{g_{s}^{3}{G^{3}}}\right\rangle m_{s}^{2}}}{{24{\pi^{2}}}}+\frac{4}{9}\left\langle{{\alpha_{s}}GG}\right\rangle{m_{s}}\left\langle{\bar{s}s}\right\rangle$
$-\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle{m_{s}}\left\langle{\bar{s}s}\right\rangle{\alpha_{s}}}}{{12\pi}}-\frac{{16\pi m_{s}^{2}{{\left\langle{\bar{s}s}\right\rangle}^{2}}{\alpha_{s}}}}{9}-\frac{4}{3}\pi\left\langle{\bar{s}s}\right\rangle\left\langle{{g_{s}}\bar{s}\sigma Gs}\right\rangle{\alpha_{s}})\,,$
$\widetilde{\Pi}^{\alpha\beta}_{1^{-+}}{\left({M_{B}^{2},{s_{0}}}\right)}$ $=\int_{4m_{s}^{2}}^{{s_{0}}}{(\frac{{{s^{3}}{\alpha_{s}}}}{{15{\pi^{3}}}}}-\frac{{4m_{s}^{2}{s^{2}}{\alpha_{s}}}}{{3{\pi^{3}}}}+s(\frac{{19\left\langle{{\alpha_{s}}GG}\right\rangle{\alpha_{s}}}}{{216{\pi^{3}}}}+\frac{{32{m_{s}}\left\langle{\bar{s}s}\right\rangle{\alpha_{s}}}}{{9\pi}})+\frac{{2\left\langle{{\alpha_{s}}GG}\right\rangle m_{s}^{2}}}{{3{\pi^{2}}}}$ (72)
$-\frac{{4{m_{s}}\left\langle{{g_{s}}\bar{s}\sigma Gs}\right\rangle{\alpha_{s}}}}{\pi})\times{e^{-s/M_{B}^{2}}}ds+(-\frac{{\left\langle{g_{s}^{3}{G^{3}}}\right\rangle m_{s}^{2}}}{{2{\pi^{2}}}}-\frac{{16}}{9}\left\langle{{\alpha_{s}}GG}\right\rangle{m_{s}}\left\langle{\bar{s}s}\right\rangle$
$-\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle{m_{s}}\left\langle{\bar{s}s}\right\rangle{\alpha_{s}}}}{{3\pi}}-\frac{{64\pi m_{s}^{2}{{\left\langle{\bar{s}s}\right\rangle}^{2}}{\alpha_{s}}}}{9}+\frac{{16}}{3}\pi\left\langle{\bar{s}s}\right\rangle\left\langle{{g_{s}}\bar{s}\sigma Gs}\right\rangle{\alpha_{s}})\,,$
$\Pi^{\alpha\beta}_{1^{-+}}{\left({M_{B}^{2},{s_{0}}}\right)}$ $=\int_{4m_{s}^{2}}^{{s_{0}}}{(\frac{{{s^{3}}{\alpha_{s}}}}{{60{\pi^{3}}}}}-\frac{{m_{s}^{2}{s^{2}}{\alpha_{s}}}}{{3{\pi^{3}}}}+s(\frac{{19\left\langle{{\alpha_{s}}GG}\right\rangle{\alpha_{s}}}}{{864{\pi^{3}}}}+\frac{{8{m_{s}}\left\langle{\bar{s}s}\right\rangle{\alpha_{s}}}}{{9\pi}})+\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle m_{s}^{2}}}{{6{\pi^{2}}}}$ (73)
$-\frac{{{m_{s}}\left\langle{{g_{s}}\bar{s}\sigma Gs}\right\rangle{\alpha_{s}}}}{\pi})\times{e^{-s/M_{B}^{2}}}ds+(-\frac{{\left\langle{g_{s}^{3}{G^{3}}}\right\rangle m_{s}^{2}}}{{8{\pi^{2}}}}-\frac{4}{9}\left\langle{{\alpha_{s}}GG}\right\rangle{m_{s}}\left\langle{\bar{s}s}\right\rangle$
$-\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle{m_{s}}\left\langle{\bar{s}s}\right\rangle{\alpha_{s}}}}{{12\pi}}-\frac{{16\pi m_{s}^{2}{{\left\langle{\bar{s}s}\right\rangle}^{2}}{\alpha_{s}}}}{9}+\frac{4}{3}\pi\left\langle{\bar{s}s}\right\rangle\left\langle{{g_{s}}\bar{s}\sigma Gs}\right\rangle{\alpha_{s}})\,,$
$\widetilde{\Pi}^{\alpha\beta}_{1^{++}}{\left({M_{B}^{2},{s_{0}}}\right)}$ $=\int_{4m_{s}^{2}}^{{s_{0}}}{(\frac{{{s^{3}}{\alpha_{s}}}}{{15{\pi^{3}}}}}-\frac{{4m_{s}^{2}{s^{2}}{\alpha_{s}}}}{{3{\pi^{3}}}}+s(\frac{{19\left\langle{{\alpha_{s}}GG}\right\rangle{\alpha_{s}}}}{{216{\pi^{3}}}}+\frac{{32{m_{s}}\left\langle{\bar{s}s}\right\rangle{\alpha_{s}}}}{{9\pi}})-\frac{{2\left\langle{{\alpha_{s}}GG}\right\rangle m_{s}^{2}}}{{3{\pi^{2}}}}$ (74)
$+\frac{{4{m_{s}}\left\langle{{g_{s}}\bar{s}\sigma Gs}\right\rangle{\alpha_{s}}}}{\pi})\times{e^{-s/M_{B}^{2}}}ds+(-\frac{{\left\langle{g_{s}^{3}{G^{3}}}\right\rangle m_{s}^{2}}}{{6{\pi^{2}}}}+\frac{{16}}{9}\left\langle{{\alpha_{s}}GG}\right\rangle{m_{s}}\left\langle{\bar{s}s}\right\rangle$
$-\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle{m_{s}}\left\langle{\bar{s}s}\right\rangle{\alpha_{s}}}}{{3\pi}}-\frac{{64\pi m_{s}^{2}{{\left\langle{\bar{s}s}\right\rangle}^{2}}{\alpha_{s}}}}{9}-\frac{{16}}{3}\pi\left\langle{\bar{s}s}\right\rangle\left\langle{{g_{s}}\bar{s}\sigma Gs}\right\rangle{\alpha_{s}})\,,$
$\Pi^{\alpha_{1}\beta_{1},\alpha_{2}\beta_{2}}_{2^{++}}{\left({M_{B}^{2},{s_{0}}}\right)}$ $=\int_{4m_{s}^{2}}^{{s_{0}}}{(\frac{{2{s^{3}}{\alpha_{s}}}}{{15{\pi^{3}}}}}-\frac{{22m_{s}^{2}{s^{2}}{\alpha_{s}}}}{{5{\pi^{3}}}}+s(\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle}}{{3{\pi^{2}}}}-\frac{{89\left\langle{{\alpha_{s}}GG}\right\rangle{\alpha_{s}}}}{{144{\pi^{3}}}}+\frac{{80{m_{s}}\left\langle{\bar{s}s}\right\rangle{\alpha_{s}}}}{{3\pi}})$ (75)
$-\frac{{2\left\langle{{\alpha_{s}}GG}\right\rangle m_{s}^{2}}}{{{\pi^{2}}}}-\frac{{12{m_{s}}\left\langle{{g_{s}}\bar{s}\sigma Gs}\right\rangle{\alpha_{s}}}}{\pi})\times{e^{-s/M_{B}^{2}}}ds+(-\frac{{{{\left\langle{{\alpha_{s}}GG}\right\rangle}^{2}}}}{{288{\pi^{2}}}}-\frac{{3\left\langle{g_{s}^{3}{G^{3}}}\right\rangle m_{s}^{2}}}{{{\pi^{2}}}}$
$+\frac{8}{3}\left\langle{{\alpha_{s}}GG}\right\rangle{m_{s}}\left\langle{\bar{s}s}\right\rangle+\frac{{2\left\langle{{\alpha_{s}}GG}\right\rangle{m_{s}}\left\langle{\bar{s}s}\right\rangle{\alpha_{s}}}}{\pi}-\frac{{176}}{3}\pi\left\langle{\bar{s}s}\right\rangle\left\langle{{g_{s}}\bar{s}\sigma Gs}\right\rangle{\alpha_{s}})\,,$
$\widetilde{\Pi}^{\alpha_{1}\beta_{1},\alpha_{2}\beta_{2}}_{2^{-+}}{\left({M_{B}^{2},{s_{0}}}\right)}$ $=\int_{4m_{s}^{2}}^{{s_{0}}}{(\frac{{8{s^{3}}{\alpha_{s}}}}{{15{\pi^{3}}}}}-\frac{{88m_{s}^{2}{s^{2}}{\alpha_{s}}}}{{5{\pi^{3}}}}+s(-\frac{{4\left\langle{{\alpha_{s}}GG}\right\rangle}}{{3{\pi^{2}}}}-\frac{{17\left\langle{{\alpha_{s}}GG}\right\rangle{\alpha_{s}}}}{{36{\pi^{3}}}}$ (76)
$+\frac{{320{m_{s}}\left\langle{\bar{s}s}\right\rangle{\alpha_{s}}}}{{3\pi}})+\frac{{8\left\langle{{\alpha_{s}}GG}\right\rangle m_{s}^{2}}}{{{\pi^{2}}}}+\frac{{48{m_{s}}\left\langle{{g_{s}}\bar{s}\sigma Gs}\right\rangle{\alpha_{s}}}}{\pi})\times{e^{-s/M_{B}^{2}}}ds$
$+(\frac{{{{\left\langle{{\alpha_{s}}GG}\right\rangle}^{2}}}}{{72{\pi^{2}}}}+\frac{{8\left\langle{g_{s}^{3}{G^{3}}}\right\rangle m_{s}^{2}}}{{{\pi^{2}}}}-\frac{{32}}{3}\left\langle{{\alpha_{s}}GG}\right\rangle{m_{s}}\left\langle{\bar{s}s}\right\rangle-\frac{{1088}}{3}\pi\left\langle{\bar{s}s}\right\rangle\left\langle{{g_{s}}\bar{s}\sigma Gs}\right\rangle{\alpha_{s}})\,,$
$\Pi^{\alpha_{1}\beta_{1},\alpha_{2}\beta_{2}}_{2^{-+}}{\left({M_{B}^{2},{s_{0}}}\right)}$ $=\int_{4m_{s}^{2}}^{{s_{0}}}{(\frac{{2{s^{3}}{\alpha_{s}}}}{{15{\pi^{3}}}}}-\frac{{2m_{s}^{2}{s^{2}}{\alpha_{s}}}}{{5{\pi^{3}}}}+s(\frac{{\left\langle{{\alpha_{s}}GG}\right\rangle}}{{3{\pi^{2}}}}-\frac{{89\left\langle{{\alpha_{s}}GG}\right\rangle{\alpha_{s}}}}{{144{\pi^{3}}}}-\frac{{16{m_{s}}\left\langle{\bar{s}s}\right\rangle{\alpha_{s}}}}{\pi})$ (77)
$+\frac{{2\left\langle{{\alpha_{s}}GG}\right\rangle m_{s}^{2}}}{{{\pi^{2}}}}+\frac{{12{m_{s}}\left\langle{{g_{s}}\bar{s}\sigma Gs}\right\rangle{\alpha_{s}}}}{\pi})\times{e^{-s/M_{B}^{2}}}ds+(-\frac{{{{\left\langle{{\alpha_{s}}GG}\right\rangle}^{2}}}}{{288{\pi^{2}}}}-\frac{{3\left\langle{g_{s}^{3}{G^{3}}}\right\rangle m_{s}^{2}}}{{{\pi^{2}}}}$
$-8\left\langle{{\alpha_{s}}GG}\right\rangle{m_{s}}\left\langle{\bar{s}s}\right\rangle+\frac{{2\left\langle{{\alpha_{s}}GG}\right\rangle{m_{s}}\left\langle{\bar{s}s}\right\rangle{\alpha_{s}}}}{\pi}+\frac{{176}}{3}\pi\left\langle{\bar{s}s}\right\rangle\left\langle{{g_{s}}\bar{s}\sigma Gs}\right\rangle{\alpha_{s}})\,,$
$\widetilde{\Pi}^{\alpha_{1}\beta_{1},\alpha_{2}\beta_{2}}_{2^{++}}{\left({M_{B}^{2},{s_{0}}}\right)}$ $=\int_{4m_{s}^{2}}^{{s_{0}}}{(\frac{{8{s^{3}}{\alpha_{s}}}}{{15{\pi^{3}}}}}-\frac{{8m_{s}^{2}{s^{2}}{\alpha_{s}}}}{{5{\pi^{3}}}}+s(-\frac{{4\left\langle{{\alpha_{s}}GG}\right\rangle}}{{3{\pi^{2}}}}-\frac{{17\left\langle{{\alpha_{s}}GG}\right\rangle{\alpha_{s}}}}{{36{\pi^{3}}}}$ (78)
$-\frac{{64{m_{s}}\left\langle{\bar{s}s}\right\rangle{\alpha_{s}}}}{\pi})-\frac{{8\left\langle{{\alpha_{s}}GG}\right\rangle m_{s}^{2}}}{{{\pi^{2}}}}-\frac{{48{m_{s}}\left\langle{{g_{s}}\bar{s}\sigma Gs}\right\rangle{\alpha_{s}}}}{\pi})\times{e^{-s/M_{B}^{2}}}ds$
$+(\frac{{{{\left\langle{{\alpha_{s}}GG}\right\rangle}^{2}}}}{{72{\pi^{2}}}}+\frac{{16\left\langle{g_{s}^{3}{G^{3}}}\right\rangle m_{s}^{2}}}{{{\pi^{2}}}}+32\left\langle{{\alpha_{s}}GG}\right\rangle{m_{s}}\left\langle{\bar{s}s}\right\rangle+\frac{{1088}}{3}\pi\left\langle{\bar{s}s}\right\rangle\left\langle{{g_{s}}\bar{s}\sigma Gs}\right\rangle{\alpha_{s}})\,.$